### v1

# NoteLLM-2: Multimodal Large Representation Models for Recommendation

Abstract
Large Language Models (LLMs) have demonstrated exceptional text understanding. Existing works explore their application in text embedding tasks. However, there are few works utilizing LLMs to assist multimodal representation tasks. In this work, we investigate the potential of LLMs to enhance multimodal representation in multimodal item-to-item (I2I) recommendations. One feasible method is the transfer of Multimodal Large Language Models (MLLMs) for representation tasks. However, pre-training MLLMs usually requires collecting high-quality, web-scale multimodal data, resulting in complex training procedures and high costs. This leads the community to rely heavily on open-source MLLMs, hindering customized training for representation scenarios. Therefore, we aim to design an end-to-end training method that customizes the integration of any existing LLMs and vision encoders to construct efficient multimodal representation models. Preliminary experiments show that fine-tuned LLMs in this end-to-end method tend to overlook image content. To overcome this challenge, we propose a novel training framework, NoteLLM-2, specifically designed for multimodal representation. We propose two ways to enhance the focus on visual information. The first method is based on the prompt viewpoint, which separates multimodal content into visual content and textual content. NoteLLM-2 adopts the multimodal In-Content Learning method to teach LLMs to focus on both modalities and aggregate key information. The second method is from the model architecture, utilizing a late fusion mechanism to directly fuse visual information into textual information. Extensive experiments have been conducted to validate the effectiveness of our method.
1 Introduction
Figure 1: Two strategies for using LLMs to enhance multimodal representations.
With the development of the internet, most online platforms are full of multimodal information to drive user engagement [51, 50, 53]. Multimodal recommendation is a necessary service for these platforms [30], which recommends items based on users’ interest on item multimodal information. These techniques mainly rely on multimodal representations, extracting raw multimodal content into dense embeddings [61], and then measuring similarity through these multimodal embeddings. The enhancement of multimodal representations can fundamentally improve numerous downstream tasks, such as retrieval [31], clustering [34], and item modeling [60, 51].
Many existing works have investigated multimodal representations, such as CLIP [37], METER [14] and Siglip [58]. Although great success has been achieved in these domains, the number of parameters in most multimodal representation models still possesses the large potential for scaling [45, 46]. However, existing works [45, 46] primarily focus on scaling vision models to improve multimodal representation, overlooking the textual branch. At the same time, due to the lack of sufficient pure text-oriented pre-training [37, 45, 46, 58], the text branch in multimodal representation models often exhibits suboptimal understanding of text information [6, 59]. Therefore, it is necessary to enhance text understanding in multimodal representations models.
Recently, Large Language Models (LLMs) have shown impressive capabilities in text understanding and generating [18, 7, 2, 47]. Several works show the superior performance of LLMs in generating textual embeddings [34, 60, 19, 31, 4]. However, the use of LLMs for enhancing text understanding in multimodal representations is underexplored. As shown in Figure 1, one feasible method is to adapt Multimodal Large Language Models (MLLMs) [29, 25, 2, 22] for representation tasks. Nevertheless, pre-training MLLMs requires collecting high-quality, web-scale multimodal data, resulting in complex training procedures and high costs [25, 2]. This leads the community to rely heavily on open-source MLLMs, hindering customized training using the most advanced LLMs and vision encoders for representation scenarios. Therefore, we aim to design an end-to-end training method that customizes the integration of any existing LLMs and vision encoders to construct efficient multimodal representation models. This end-to-end method is practical when the target LLMs lack pre-trained visual components [11, 18, 1], as it only needs target domain data, eliminating pre-training data collection. Besides, we can optimize the customized model structure for representation tasks to achieve better efficiency while ensuring performance. We call these models for enhancing multimodal representation with LLMs as Multimodal Large Representation Models (MLRMs).
In this work, we explore MLRMs in item-to-item (I2I) multimodal recommendation scenarios. We conduct fine-tuned preliminary experiments to examine our goal. Our findings reveal that representations of naive fine-tuned end-to-end MLRMs are biased towards textual information, ignoring visual content. Visual information is mainly used to amplify the key element in multimodal content, rather than being directly aggregated by representations. To solve this challenge, we propose a novel training framework, called NoteLLM-2, for multimodal representation tasks. We design two methods to mitigate visual modality neglect. The first method is multimodal In-Context Learning (mICL). This method separates multimodal content into visual and textual components, subsequently compressing the content into two modality-compressed words. Each modality-compressed word conducts in-batch contrastive learning with its corresponding modality-compressed word. During inference, mICL can aggregate multimodal information in an In-Context Learning (ICL) way [12, 48]. The second method is the late fusion mechanism. Most MLLMs adopt early fusion, which maps visual features into the textual feature space of LLMs and interact visual features with text embeddings through LLMs [29, 25, 2, 22]. Even though early fusion can deepen the multimodal information interaction, it can easily lead to the loss of pure visual information during the multimodal fusion conducted by LLMs [25, 62]. We incorporate late fusion based on early fusion, retaining more visual information by delaying the fusion process. Specifically, we adopt a multimodal gate fusion mechanism [8], which takes visual representations from the visual encoder and representations from the LLM to generate final multimodal representations. We conduct extensive experiments to demonstrate our NoteLLM-2 effectiveness. Our paper makes the following contributions:
  • To the best of our knowledge, we are the first to explore the multimodal representation assisted by LLMs. We propose an end-to-end training method that customizes the integration of any existing LLMs and vision encoders to form efficient MLRMs.
  • We reveal that fine-tuning unaligned MLRMs for representation tasks can cause the modality imbalance, overlooking visual information. To overcome this problem, we design a novel framework, NoteLLM-2, which contains two mechanisms, mICL and late fusion.
  • Extensive experiments have demonstrated the effectiveness of our proposed framework for multimodal recommendation representation.
Our paper is organized as follows: We begin with a preliminary investigation of MLRMs in multimodal I2I recommendation scenarios and analyze the results. Then, inspired by this analysis, we design our NoteLLM-2. And we conduct experiments to demonstrate the effectiveness of our NoteLLM-2. Besides, related works and additional details are presented in Appendix A.
2 Preliminary Investigation
2.1 Preliminaries
Problem Statement. In our scenarios, every item represents a note, which is generated by users and expresses their life experiences. We show some notes in Appendix H. We assume $\mathcal{N}=\{n_{1},n_{2},...,n_{m}\}$ as note pool, where $m$ is the number of notes. Each note contains textual information, like the title, topic, and content, as well as the image. We denote the $i$-th note as $n_{i}=(t_{i},tp_{i},ct_{i},v_{i})$, where $t_{i}$, $tp_{i}$, $ct_{i}$, $v_{i}$ mean the title, the topic, the content and the image respectively. Given a query note $n_{i}$, the I2I recommendation system ranks the note pool $\mathcal{N}\backslash\{n_{i}\}$ according to the multimodal content similarity to the given note. The objective of this system is to prioritize the ranking of the related target note.
Dataset Construction. Following [60], we utilize the co-occurrence mechanism to build related note pairs based on user behaviors without human annotations. This mechanism assumes that notes frequently read together are likely related. Then, we count the co-occurrence score in which users viewed note $n_{A}$ and subsequently clicked on note $n_{B}$ as follows:
$s_{n_{A}\rightarrow n_{B}}=\sum_{u\in U_{n_{A}\rightarrow n_{B}}}\frac{1}{N_{u}},$ (1)
where $s_{n_{A}\rightarrow n_{B}}$ means the co-occurrence score from note $n_{A}$ to note $n_{B}$, $U_{n_{A}\rightarrow n_{B}}$ is the set of users who viewed note $n_{A}$ and subsequently clicked on note $n_{B}$, and $N_{u}$ denotes the quantity of the note set clicked by the user $u$ in the user behavior data. After calculating the co-occurrence score for all note pairs, we construct the set of co-occurrence scores $\mathcal{S}_{n_{i}}$ from note $n_{i}$ to all other notes. Specifically, $\mathcal{S}_{n_{i}}$ is defined as $\{{s_{n_{i}\rightarrow n_{j}}}|1\leq j\leq m,i\neq j\}$. Next, we remove outlier from $\mathcal{S}_{n_{i}}$ with co-occurrence scores above $up$ or below $low$. Finally, we select the $t$ notes with the highest co-occurrence scores from the filtered set as the related notes for note $n_{i}$.
Figure 2: The frameworks of MLRMs. (a) is the basic method for representations. (b) is our representation method, which contains two easy and effective mechanisms to enhance multimodal representation ability, mICL and late fusion.
Basic Representation Method of MLRMs. Following [19], we adopt a prompt with explicit one word limitation to compress the multimodal content into one embedding. Specifically, we use json format to show note content and our prompt is the following: {mdframed} Note Compression Prompt
Note content: {’image’: <IMG>, ’title’: $t_{i}$ , ’topic’: $tp_{i}$ , ’content’: $ct_{i}$}. Compress this note into one word:" In this template, <IMG> is a placeholder, which will be replaced by vision embeddings processed from raw images $v_{i}$.
The overall framework is shown in Figure 2 (a). This framework first utilizes the vision encoder $V_{\theta}$ to extract the pre-processed image $v_{i}$ into visual features $Z_{v}\in\mathbb{R}^{L\times h_{v}}$ in the notes $n_{i}$, where $L$ is the length of visual features, $h_{v}$ is the dimension of visual features and $Z_{v}=V_{\theta}(v_{i})$. Then, the connector $C_{\pi}$ transforms the visual features $Z_{v}$ into the word embedding space of LLMs to form visual embeddings $E_{v}\in\mathbb{R}^{L_{c}\times h_{t}}$, where $L_{c}$ is the length of visual embeddings input into LLMs, $h_{t}$ is the dimension of LLMs’ hidden states and $E_{v}=C_{\pi}(Z_{v})$. After processing images, the text prompts are tokenized into discrete indexes, then form the text word embedding $E_{t}\in\mathbb{R}^{T\times h_{t}}$, where $T$ is the length of text tokens. To insert the visual embeddings into the correct position in text embeddings, we replace word embeddings in positions of the <IMG> token by the visual embeddings $E_{v}$, forming the multimodal embeddings $E_{m}\in\mathbb{R}^{(L_{c}+T-1)\times h_{t}}$. Last, we use LLMs $LLM_{\mu}$ to process the multimodal embeddings to generate last hidden states $H\in\mathbb{R}^{(L_{c}+T-1)\times h_{t}}$, where $H=LLM_{\mu}(E_{m})$. We take last embeddings of $H$ as the note representation $\boldsymbol{n}_{i}$ for note $n_{i}$. This method constrains LLMs to compress the multimodal content into one embedding in the form of the next token prediction.
However, LLMs are trained using the language modeling loss [29, 25, 3, 22], which significantly differs from representation tasks. To bridge this gap, we employ contrastive learning, a prevalent technique for training embeddings [34, 60, 19, 31]. Specifically, each minibatch contains $B$ related note pairs $(n_{i},n_{i}^{+})$, resulting in a total of $2B$ notes per minibatch. Following [36], we use gradient descent to minimize the contrastive loss as follows:
$\mathcal{L}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\boldsymbol{n}_{i},\boldsymbol{n}_{i}^{+})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\boldsymbol{n}_{i},\boldsymbol{n}_{j})\cdot e^{\tau}}},$ (2)
where $\tau$ means the learnable temperature and $sim(a,b)=a^{\top}b/(\|a\|\|b\|)$. $\mathcal{L}(\pi,\mu)$ means that we only update the connector $C_{\pi}$ and LLMs $LLM_{\mu}$, while keep the vision encoder $V_{\theta}$ frozen, aiming for the larger batch size and better performance [22].
2.2 Datasets and Experimental Settings
Datasets. We obtain a real-world multimodal I2I dataset from our platform. We create the training dataset by randomly choosing related note pairs from user behaviour data gathered over two weeks. One-tenth of pairs in this training dataset is used for the validation set. Next, we randomly select notes from the upcoming week to form the note pool of testing set, excluding any notes that are already in the training dataset. Then, we count the related note pairs in the note pool of testing set. The training dataset contains $1.5M$ notes and $1.1M$ note pairs. The testing dataset contains $0.5M$ notes and $21K$ note pairs. We provide detailed statistics of the training and testing dataset in Appendix C.
To better evaluate the multimodal representation ability of MLRMs rather than relying heavily on the text modality, we collect pairs that include short notes in the test dataset. We define notes with a token length of less than $50$ as short notes, which approximately constitute $10\%$ of the total test notes. We categorize pairs with short query notes as short query pairs and those with short target notes as short target pairs. In the test dataset, the number of short query pairs is $5,620$ and the number of short target pairs is $5,582$.
Experimental settings. In related note pair construction, we set the upper bound of the co-occurrence score $up$ as $30$ and the lower bound $low$ as $0.01$. And we set $t$ as $3$. To maintain context length restrictions, we truncate the titles exceeding $20$ tokens, and truncate the contents exceeding $80$ tokens. In the fine-tuned experiments, for fair comparison, we add a linear projector to diminish the dimension of note embeddings to $64$. The batch size $B$ is set to 128128128128111We train models on 8 $\times$ 80GB Nvidia A100 GPUs and the batch size per each GPU is $16$.. Each batch contains $256$ notes. The temperature $\tau$ is initialized as $3$. We provide more training details in Appendix D.
To evaluate the representation ability of MLRMs, we rank all notes in the note pool, excluding the query note, according to the content of query note. We compute recall scores based on the position of the target note in the ranked list. We report Recall@100, Recall@1k and Recall@10k on all pairs, short query pairs and short target pairs in the test dataset.
Table 1: Representation performance of different models (%).
  Method Input All Pair Short Query Pair Short Target Pair
R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
                          Unimodal Training
  CLIP ViT-B Image 23.21 45.44 71.16 25.24 49.40 75.99 25.58 50.28 75.80
BM25 Text 55.20 72.67 82.83 35.82 48.42 59.71 35.07 49.63 63.68
RoBERTa-wwm-ext Text 67.59 85.06 93.86 41.23 58.61 76.40 41.80 60.65 80.04
Tomato Text 71.01 87.38 95.19 43.55 60.70 79.10 45.04 63.45 83.24
Tomato + NoteLLM Text 71.42 87.78 95.57 44.13 61.86 80.07 45.73 64.28 84.22
Qwen-Chat Text 73.03 88.41 97.50 45.52 62.21 79.93 46.42 64.49 83.62
                          Multimodal Training
  METER Merge-attn Multimodal 68.22 88.37 97.48 50.37 75.61 93.59 50.05 76.55 94.67
METER Co-attn Multimodal 70.90 89.75 97.88 52.70 78.14 94.65 53.17 79.74 95.65
BLIP-2 Multimodal 68.38 88.22 97.44 52.97 77.83 94.40 54.00 79.74 95.29
MTomato-Base Multimodal 71.94 88.22 96.13 44.77 63.77 83.31 45.83 66.08 87.02
MQwen-Base Multimodal 74.02 89.65 97.22 48.15 68.35 88.22 49.82 70.91 91.43
MQwen-bigG Multimodal 77.64 92.89 98.91 57.45 80.92 95.73 57.90 83.49 96.99
Qwen-VL-Chat [^2] Due to Qwen-VL-Chat having $256$ visual embeddings per image, this model is trained on 32 $\times$ 80GB Nvidia A100 GPUs and the batch size per each GPU is $4$. Multimodal 78.54 93.77 99.03 60.43 83.26 96.60 61.63 85.48 97.89
2.3 Performance of Multimodal Representation of Fine-tuned MLRMs
We conduct zero-shot experiments on several existing MLLMs in Appendix G, and find that zero-shot is insufficient to adapt MLLMs to representation tasks, performing inferiorly compared to a simple baseline, BM25 [40]. This is due to the misalignment between pre-training tasks and representation tasks [60, 19]. Hence, fine-tuning MLLMs for representation tasks is necessary.
We design three end-to-end MLRMs to examine the representation training method: MTomato-Base uses Tomato (our continually pre-trained LLM based on LLaMA 2 [47], which lacks vision perception ability) as the LLM, CLIP ViT-B [37] as the vision encoder and a randomly initialized Q-Former [25] as the connector for efficiency. MQwen-Base replaces Tomato with Qwen-Chat [2] in MTomato-Base. MQwen-bigG replaces CLIP ViT-B with ViT-bigG [17] in MQwen-Base. We set the visual embedding length of these models as $16$ for efficiency. For comparison, we choose two pre-trained MLLMs: BLIP-2 and Qwen-VL-Chat [3]. All vision encoders in these models are frozen for the larger batch size. In addition to MLRMs, we also compare other existing multimodal representation baselines. We provide details about baselines and MLRMs in Appendix E.
The results are shown in Table 1. Several interesting observations can be made. Firstly, MLRMs based on LLMs can outperform existing baselines significantly. Qwen-VL-Chat achieves a $10.78\%$ improvement in R@100 compared to the traditional method, METER Co-attn on all pairs. Secondly, the end-to-end training representation method can enhance the model’s multimodal representation ability. MQwen-bigG improves R@100 on all pairs by $6.31\%$ compared to Qwen-Chat. However, when the visual encoder is small, such as CLIP ViT-B, the enhancement in multimodal perception is not substantial. MTomato-Base only improves R@100 on all pairs by $1.31\%$ compared to Tomato, and MQwen-Base only improves R@100 on all pairs by $1.36\%$ compared to Qwen-Chat. Lastly, although MQwen-bigG is more efficient than Qwen-VL-Chat, which uses the same vision encoder and LLM (see Appendix E.3), there remains a performance gap.
Table 2: Representation performance of fine-tuned MLRMs under different modal inputs (%).
  Method Input All Pair Short Query Pair Short Target Pair
R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
                          Multimodal Training
  BLIP-2 Image 24.72 45.21 69.68 26.07 48.78 75.59 27.59 49.78 75.97
Text 56.85 75.56 88.26 32.85 47.28 64.66 34.88 51.68 72.98
Multimodal 68.38 88.22 97.44 52.97 77.83 94.40 54.00 79.74 95.29
  MTomato-Base Image 0.60 1.49 6.98 0.54 1.51 9.15 0.84 2.13 9.70
Text 71.49 87.59 95.31 43.80 61.74 79.26 44.72 63.89 83.26
Multimodal 71.94 88.22 96.13 44.77 63.77 83.31 45.83 66.08 87.02
  MQwen-Base Image 1.92 6.59 23.43 2.86 8.54 28.83 3.05 9.38 29.72
Text 73.14 88.35 95.34 44.50 61.59 78.78 46.12 64.41 83.97
Multimodal 74.02 89.65 97.22 48.15 68.35 88.22 49.82 70.91 91.43
  MQwen-bigG Image 18.32 38.50 64.56 19.76 41.56 71.32 21.30 42.72 71.54
Text 70.63 86.09 93.81 39.94 55.77 72.17 43.03 60.02 79.77
Multimodal 77.64 92.89 98.91 57.45 80.92 95.73 57.90 83.49 96.99
  Qwen-VL-Chat Image 24.67 45.51 70.08 25.94 48.32 75.32 26.94 49.59 75.67
Text 71.44 86.78 94.43 42.55 58.59 75.86 43.97 61.11 80.17
Multimodal 78.54 93.77 99.03 60.43 83.26 96.60 61.63 85.48 97.89
We also evaluate all fine-tuned MLRMs under different modal inputs, as shown in Table 2. Firstly, we observe that MTomato-Base and MQwen-Base hardly represent the image content when the input consists solely of images, which shows ignoring visual information. Owing to its more robust vision encoder, MQwen-bigG exhibits superior vision representation capabilities. However, its performance still falls short when compared to BLIP-2 and Qwen-VL-Chat, both of which have undergone multimodal pre-training.
2.4 Exploring the Bias of Fine-tuned MLRMs
To further mine the bias in fine-tuned MLRMs, we explore the attention patterns of two modality tokens in different LLMs’ layers. Following [48], we adopt saliency scores to analyze the interaction between the compressed word and two modality tokens in the attention matrices. We randomly select $10,000$ notes in batches from the training dataset to count the saliency scores, and compute the saliency scores as follows:
$I_{l}=\sum_{h}\left|A^{\top}_{h,l}\frac{\partial\mathcal{L}}{\partial A_{h,l}}\right|,$ (3)
where $A_{h,l}$ means the attention matrix value of the $h$-th attention head in the $l$-th layer and $\mathcal{L}$ is the contrastive loss. We sum all attention heads to obtain the saliency matrix $I_{l}$ of $l$-th layer. $I_{l}(i,j)$ means the significance of the information flow from the $j$-th word to the $i$-th word. To analyze different modality importances for the compressed word, we separate $I_{l}$ into three parts as flowing: $S_{v}$, the mean significance of information flow from the vision embeddings $E_{v}$ to the final compressed position $c$. $S_{t}$, the mean significance of information flow from the textual embeddings $E_{t}$ to the final compressed position $c$. $S_{o}$, the mean significance of information flow amongst all words, excluding influences represented by $S_{v}$ and $S_{t}$. We provide a more clear formal definition in Appendix F.1.
The results of three parts from the saliency matrix are shown in Figure 3. MLRMs trained end-to-end without pre-training exhibit a low $S_{v}$ and a high $S_{o}$ in shallow layers, while pre-trained MLLMs, BLIP-2 and Qwen-VL-Chat (see Appendix F.2), have more balanced $S_{v}$ and $S_{o}$ in shallow layers. This indicates that end-to-end trained MLRMs primarily utilize the information gain from multimodal data to enhance the information flow in the context. Images in these models primarily assist the information flow in identifying crucial information [44, 43], rather than being directly aggregated by representations. Besides, text information flow significantly surpasses visual information flow in end-to-end trained MLRMs.
((a)) BLIP-2
((b)) MTomato-Base
((c)) MQwen-Base
((d)) MQwen-bigG
Figure 3: Relative sizes of $S_{v}$, $S_{t}$ and $S_{o}$ in different layers of different MLRMs.
3 Methodology
From our investigation, MLRMs trained end-to-end without pre-training are suboptimal, primarily due to the disregard of visual information after processing by LLMs. Therefore, we aim to design mechanisms that focus more on visual signals. We propose a novel training framework, NoteLLM-2, which contains two methods from different perspectives. The first one is from the prompt viewpoint, called mICL. This method alters the prompt to adjust the attention patterns towards visual information. The second one is from the model architecture perspective, which combines the late fusion with the visual prompt. This enhances the impact of the visual information flow on the final representation by delaying the fusion of visual information. The improved overall framework is shown in Figure 2 (b).
Specifically, given note $n_{i}$, the mICL mechanism does not attempt compress the multimodal information into one compressed word. Instead, we separate the multimodal note into two single-modality notes. We then adopt a similar ICL way [12, 48] to aggregate the multimodal information. We reformulate our note compression prompt as follows: {mdframed} mICL Note Compression Prompt.
Note content: {’image’: <IMG>}, Compress this note into one word:"<IMG_EMB>". Note content: {’title’: $t_{i}$ , ’topic’: $tp_{i}$ , ’content’: $ct_{i}$}. Compress this note into one word:" In this template, <IMG_EMB> is a special token.
After processing the multimodal embeddings with LLMs, we choose relevant hidden states to represent notes. Following [60, 19], we take the previous hidden state of the token <IMG_EMB> as the visual note representation, denoted as $\boldsymbol{n}_{i}^{v}$. Due to the causal attention, $\boldsymbol{n}_{i}^{v}$ only contains the note image information. We also take the last hidden state as the multimodal note representation, denoted as $\boldsymbol{n}_{i}^{m}$, which can compress visual and textual information. This method separates the original single embedding into two modality note embeddings. The multimodal note representation $\boldsymbol{n}_{i}^{m}$ aggregates useful visual information in a similar manner to ICL.
The late fusion mechanism adopts original visual embedding to enhance the note embeddings. This prevents a textual bias due to the space of LLMs [62] and incorporates more original visual information. Specifically, after extracting image features by the vision encoder $V_{\theta}$, we take the visual feature, which contains the entire image information, such as the [CLS] token in CLIP ViT-B. Then, we use a linear layer to transform these features into the dimension of LLMs, denoted as $\boldsymbol{v}\in\mathbb{R}^{h_{t}}$. We adopt the same gate mechanism to fuse the original visual information into the two note embeddings.
$\boldsymbol{z}$ $=sigmoid(\boldsymbol{W}[\boldsymbol{v},\boldsymbol{n}_{i}^{v}]+\boldsymbol{b}),$ (4)
$\hat{\boldsymbol{n}_{i}^{v}}$ $=\boldsymbol{z}\odot\boldsymbol{v}+(\boldsymbol{1}-\boldsymbol{z})\odot\boldsymbol{n}_{i}^{v},$
$\boldsymbol{z}$ $=sigmoid(\boldsymbol{W}[\boldsymbol{v},\boldsymbol{n}_{i}^{m}]+\boldsymbol{b}),$ (5)
$\hat{\boldsymbol{n}_{i}^{m}}$ $=\boldsymbol{z}\odot\boldsymbol{v}+(\boldsymbol{1}-\boldsymbol{z})\odot\boldsymbol{n}_{i}^{m},$
where $\hat{\boldsymbol{n}_{i}^{v}}$ and $\hat{\boldsymbol{n}_{i}^{m}}$ means the fused note embeddings. [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] means the concatenate operation, and $\boldsymbol{W}$ and $\boldsymbol{b}$ are learnable parameters. $\odot$ is the element-wise product.
Next, we adopt the two fused embeddings to conduct contrastive learning as follows:
$\mathcal{L}^{v}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\hat{\boldsymbol{n}_{i}^{v}},\hat{\boldsymbol{n}_{i}^{v,+}})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\hat{\boldsymbol{n}_{i}^{v}},\hat{\boldsymbol{n}_{j}^{v}})\cdot e^{\tau}}},$ (6)
$\mathcal{L}^{m}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\hat{\boldsymbol{n}_{i}^{m}},\hat{\boldsymbol{n}_{i}^{m,+}})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\hat{\boldsymbol{n}_{i}^{m}},\hat{\boldsymbol{n}_{j}^{m}})\cdot e^{\tau}}},$ (7)
where $\mathcal{L}^{v}(\pi,\mu)$ is the loss from visual note embeddings and $\mathcal{L}^{m}(\pi,\mu)$ is the loss from the multimodal note embeddings. The final loss is computed as follows:
$\mathcal{L}^{f}(\pi,\mu)=\frac{\mathcal{L}^{v}(\pi,\mu)+\alpha\mathcal{L}^{m}(\pi,\mu)}{1+\alpha},$ (8)
where $\mathcal{L}^{f}(\pi,\mu)$ is the final loss and $\alpha$ is a hyperparameter. In the evaluation, we use $\hat{\boldsymbol{n}_{i}^{m}}$ as the note embeddings, which contain multimodal information.
4 Experiments
Table 3: Performance of NoteLLM-2 based on three different MLRMs (%).
  Method All Pair Short Query Pair Short Target Pair
R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
  MTomato-Base 71.94 88.22 96.13 44.77 63.77 83.31 45.83 66.08 87.02
   + mICL 74.16 90.95 98.26 51.56 75.86 93.84 52.71 78.12 95.82
   + late fusion 74.51 91.37 98.45 54.19 78.02 94.50 54.00 79.71 95.80
   + NoteLLM-2 74.61 91.50 98.47 54.87 78.81 94.69 54.52 79.75 95.46
only late fusion 74.22 91.25 98.43 53.21 77.75 94.65 53.65 79.77 95.59
  MQwen-Base 74.02 89.65 97.22 48.15 68.35 88.22 49.82 70.91 91.43
   + mICL 75.82 91.57 98.55 53.32 76.40 94.38 54.52 79.31 96.05
   + late fusion 75.92 91.49 98.28 52.74 76.03 93.65 54.06 78.89 95.53
   + NoteLLM-2 76.50 92.18 98.58 55.72 78.91 94.94 56.13 80.65 96.24
only late fusion 75.84 91.49 98.35 53.36 76.23 93.57 54.27 78.39 95.30
  MQwen-bigG 77.64 92.89 98.91 57.45 80.92 95.73 57.90 83.49 96.99
   + mICL 77.97 93.19 98.92 58.50 82.08 96.66 58.58 84.01 97.20
   + late fusion 77.43 92.92 98.78 57.47 81.77 95.89 58.50 83.99 97.16
   + NoteLLM-2 77.56 93.38 98.82 57.94 81.98 96.29 58.79 83.87 97.05
only late fusion 76.24 92.31 98.67 55.21 79.22 95.58 55.15 81.13 96.70
4.1 Performance Evaluation
We conduct experiments on all components of NoteLLM-2 based on three MLRMs to show our method effectiveness. Besides, we perform an ablation study, only late fusion, which only use late fusion to integrate image and text information, without inputting image embeddings into LLMs.
The results are presented in Table 3. Firstly, NoteLLM-2 can significantly enhance the overall performance for MTomato-Base and MQwen-Base, which have relatively small vision encoders. Additionally, NoteLLM-2 primarily enhances MQwen-bigG in terms of short pairs. Secondly, mICL enhances the performance of all models, while late fusion is more effective in models with relatively small vision encoders. Concurrently, we find that only late fusion is a straightforward and efficient fusion method. However, its simplicity could potentially be a drawback when the vision encoder is more powerful, as it may not fully and effectively interact with LLMs, leading to a loss in performance. Lastly, MQwen-bigG + NoteLLM-2 still underperforms compared to Qwen-VL-Chat, especially in shorter pairs. We attribute this to the significant difference in visual token length between MQwen-bigG ($16$) and Qwen-VL-Chat ($256$).
((a)) MTomato-Base
((b)) MQwen-Base
((c)) MQwen-bigG
Figure 4: Relative sizes of saliency scores in different layers of different MLRMs.
4.2 Saliency Scores of Enhanced MLRMs
To further understand the impact of NoteLLM-2 on MLRMs, we illustrate the difference in saliency scores of enhanced MLRMs. The saliency scores for the original fine-tuned method are $S_{v}$, $S_{t}$ and $S_{o}$. The saliency scores for the enhanced fine-tuned method are $\hat{S_{v}}$, $\hat{S_{t}}$ and $\hat{S_{o}}$. We consider the visual note compressed word as a part of the vision embeddings $E_{v}$. Our results are shown in Figure 4. The representations of all enhanced MLRMs amplify the direct focus on images while reducing the emphasis on $S_{o}$ in the shallow layers. At the same time, $S_{t}$ remains virtually unchanged. This is attributed to mICL’s use of the same compressed prompt to compress both modalities. By identifying similar compression patterns from image information, mICL can bolster the concentration of multimodal representations on images, similar to ICL [48].
Table 4: Impact of the number of visual tokens $L$ (%).
  #Visual Tokens All Pair Short Query Pair Short Target Pair
R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
  8 74.17 90.85 98.16 52.80 76.13 93.76 53.19 77.87 94.48
16 74.61 91.50 98.47 54.87 78.81 94.69 54.52 79.75 95.46
32 74.55 91.41 98.38 54.56 78.91 94.94 54.21 79.83 95.63
48 74.30 91.57 98.48 54.31 78.83 95.13 53.73 80.06 95.92
Table 5: Impact of textual loss and visual loss ratio $\alpha$ (%).
  $\alpha$ All Pair Short Query Pair Short Target Pair
R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
  1 73.94 91.11 98.38 54.13 78.16 94.67 54.06 79.00 95.38
3 74.38 91.41 98.42 54.15 77.71 94.59 54.06 79.31 95.53
9 74.61 91.50 98.47 54.87 78.81 94.69 54.52 79.75 95.46
19 74.53 91.33 98.54 54.56 78.68 95.00 53.92 80.04 95.53
4.3 Hyper-Parameter Analysis
In this section, we use MTomato-Base to conduct hyper-parameter analysis experiments.
Impact of length of visual tokens. We explore the impact of visual token length. The results are shown in Table 4. Reducing the length from $16$ to $8$ worsens performance in short pairs. There exists a tradeoff between performance and inference speed concerning the length of visual tokens.
Impact of text and visual loss ratio. We vary the ratio of text to visual loss $\alpha$, as shown in Table 5. When the ratio is small, the performance is slightly worse. However, as we increase the ratio, we find that our method is insensitive to changes in the text to visual loss ratio.
5 Conclusion
This paper explores using LLMs to improve textual comprehension in multimodal representation tasks under I2I recommendation scenarios. We design an end-to-end training method that can customize the integration of any existing LLMs and vision encoders. This approach reduces reliance on open-source MLLMs, which require costly multimodal pre-training. To prevent the problem of ignoring visual information in end-to-end training, we propose NoteLLM-2, which contains mICL and late fusion. The effectiveness of this approach is confirmed through extensive experiments.
Appendix
Appendix A Related Work
A.1 Multimodal Representation
With the advancement of pre-training techniques, numerous studies utilize web-scale image-text data to pre-train models for a fundamental understanding of multimodal information. There are two main pre-training techniques. The first technique adopts generative pre-training objectives, following masked language/image modeling [14, 49, 27] or image captioning [56, 26]. The second technique uses contrastive pre-training objectives, such as contrastive learning [37, 58, 45, 46, 56, 49] and image-text matching tasks [14, 27, 26]. These models demonstrate a powerful zero-shot ability in many downstream tasks. However, the text comprehension capabilities of these models are unsatisfactory [6, 59], particularly in the era of LLMs. In this work, we explore to use LLMs to enhance the multimodal representations.
A.2 Large Language Models
LLMs have attracted extensive attention due to their remarkable text understanding and generation capability [18, 7, 2, 47]. Therefore, numerous text applications have been developed based on LLMs [31, 60, 34, 52, 22]. To understand more modality information, LLMs are extended into MLLMs [3, 5, 55, 33, 24]. Most MLLMs employ specialized modality encoders to capture non-textual information [42, 3, 5]. This information is then mapped into textual LLM space using connectors and subsequently processed by LLMs. Due to the significant understanding ability of LLMs, MLLMs can generate answers according to textual instruction and multimodal information [15, 13, 3, 5]. While existing studies primarily focus on the generation abilities of LLMs, some explore the text representation ability of LLMs [34, 41, 60, 19, 31, 4], demonstrating significant potential to enhance performance for numerous downstream tasks. However, no studies have explored the multimodal representation ability with the assistance of LLMs. Therefore, we aim to investigate the multimodal representation ability enhanced by LLMs in recommendation scenarios.
A.3 Multimodal I2I Recommendation
I2I recommendation generates a ranked list of items from a vast pool based on a specific target item, which is an important technique for recall stages in recommendation [54, 64, 60]. I2I recommendation employs either a pre-constructed I2I index [54] or the approximate k-nearest neighbor method [21] to retrieve relevant items online. Some studies have investigated multimodal content-based I2I recommendations [57, 10, 20, 63, 16]. For example, CommerceMM [57] adopts multimodal pre-training tasks used in METER [14] and omni-retrieval tasks to train commerce multimodal representation. MAPS [10] proposes a multimodal attention framework to measure product similarity. However, these works adopt small-scale pre-trained visual-linguistic models as backbones, which have huge potential for scaling up. As shown in [60, 19, 4], LLMs can outperform smaller models in textual representation tasks significantly. Therefore, in this work, we explore empowering the multimodal representations with LLMs in multimodal I2I recommendation.
Appendix B Limitations and Future Works
More modalities. In this work, our focus is primarily on the processing of multimodal situations that involve the input of a single image and corresponding text information. This is the most simplified situation within our research scenarios. In fact, the scope of our representation method can extend to include audio and video data [42, 32, 28]. This direction presents an opportunity for further exploration and expansion into more complex modalities.
General scenarios. Our MLRMs mainly focus on the multimodal representation ability in recommendation scenarios. These scenarios determine item similarity by analyzing the interaction between items and users. However, this is not a universally applicable scenario. The current focus of our MLRMs is somewhat specialized and may not cover all potential use cases. Given the potential of MLRMs, it is promising to consider broadening their multimodal representation ability into more general scenarios [35].
Larger and stronger base models. Due to the rapid development of LLMs and vision encoders, there are an extensive range of larger and more powerful base models available in the market [58, 2, 11]. However, due to resource limitations and the demand for rapid recommendations, we have not yet experimented with all of these models. There is still a vast scope to investigate more advanced models, which could potentially boost performance even further.
Appendix C Dataset Details
Table 6: Detailed statistics of training and testing dataset.
                                  training dataset
  # notes 1,487,613 # note pairs 1,130,906
avg. # words per title 12.38 avg. # topic per note 5.26
avg. # words per topic 4.53 avg. # words per content 133.20
                                  testing dataset
  # notes 534,144 # note pairs 21,262
avg. # words per title 11.68 avg. # topic per note 4.58
avg. # words per topic 4.50 avg. # words per content 116.00
Appendix D Training Details
Table 7: Detailed training hyperparameters for training MLRMs.
  Configuration Hyperparameter
  Optimizer AdamW
$\beta_{1}$ $0.9$
$\beta_{2}$ $0.999$
$epsilon$ $1e^{-8}$
Max gradient norm $1.0$
Weight decay $1e^{-3}$
Peak learning rate $3e^{-6}$
Warmup Ratio $0.1$
Learning rate scheduler linear decay
Numerical precision bf16
Global batch size $128$
Training steps $7,956$
For all models with LLMs, we list the detailed training hyperparameter in Table 7. Besides, we use DeepSpeed [38], and Zero Redundancy Optimizer (ZeRO) [39] Stage 3 to train our models.
Appendix E Model Details
E.1 Baseline Details
To evaluate the effectiveness of MLRMs, we compare them with several baselines. (1) CLIP ViT-B [37] is trained using only image input. (2) BM25 [40] is a basic baseline, which counts the relevance between the query and documents based on every word in the query. (3) RoBERTa-wwm-ext [9] is a classic traditional Chinese BERT-like [23] text encoder with pure text input. (4) Tomato is our continually pre-trained LLM that is based on LLaMA 2 [47] and utilizes data from our platform for training. (5) NoteLLM [60] unifies the embedding generation and topic generation into one task. (6) METER Merge-attn and METER Co-attn [14] leverage best training strategy to train vision-and-language transformers. METER Merge-attn adopts the merged attention as the fusion module, while METER Co-attn uses co-attention as the fusion module. All baselines are trained on our training dataset.
E.2 End-to-End MLRM Details
We provide the details in Table 8.
Table 8: Model settings of end-to-end MLRMs.
  Configuration MTomato-Base MQwen-Base MQwen-bigG
  Vision encoder $V_{\theta}$ init. CLIP ViT-B CLIP ViT-B OpenCLIP ViT-bigG
Connector $C_{\pi}$ init. random Q-Former random Q-Former random Q-Former
LLM $LLM_{\mu}$ init. Tomato Qwen-Chat Qwen-Chat
  # Vision encoder layers $12$ $12$ $48$
# Vision encoder attention heads $12$ $12$ $16$
Vision encoder sequence length $L$ $196$ $196$ $256$
Vision encoder hidden size $h_{v}$ $768$ $768$ $1,664$
Vision encoder intermediate size $3,072$ $3,072$ $8,192$
  # Q-Former layers $6$ $6$ $6$
# Q-Former attention heads $12$ $12$ $12$
Learnable query numbers $L_{c}$ $16$ $16$ $16$
Q-Former cross attention frequency $2$ $2$ $2$
Q-Former cross attention hidden size $768$ $768$ $1,664$
Q-Former hidden size $768$ $768$ $1,536$
Q-Former intermediate size $3,072$ $3,072$ $3,072$
  # LLM layers $32$ $32$ $32$
# LLM attention heads $32$ $32$ $32$
Vocab size $49,216$ $151,936$ $151,936$
LLM hidden size $h_{t}$ $4,096$ $4,096$ $4,096$
LLM intermediate size $11,008$ $22,016$ $22,016$
E.3 Comparison of MLRMs
We provide the comparison in Table 9.
Table 9: Comparison of different LLMs for representation tasks. $L_{c}$ means the number of vision embeddings for each image. The inference speed test is conducted on an A100 GPU, with an inference batch size set at $50$. All notes in this test are the longest note in the dataset. The speed is quantified as the number of longest notes inferred per second.
  Method Vision Encoder LLM VL Connector # Total Par. $L_{c}$ Inf. Speed
  Tomato - Tomato (6.7B) - 6.7B - 59.0
Qwen-Chat - Qwen-Chat (7.1B) - 7.1B - 57.0
  BLIP-2 ViT-g (1.0B) OPT (6.7B) Q-Former (105.1M) 7.8B 32 44.2
LLaVA-1.5 ViT-L (303.5M) Vicuna (6.7B) MLP (21.0M) 7.1B 576 12.1
Qwen-VL-Chat ViT-bigG (1.8B) Qwen-Chat (7.1B) Resamplers (76.1M) 9.0B 256 11.5
MTomato-Base ViT-B (85.8M) Tomato (6.7B) Q-Former (49.6M) 6.8B 16 51.5
    + NoteLLM-2 ViT-B (85.8M) Tomato (6.7B) Q-Former (49.6M) 6.9B 16 48.2
MQwen-Base ViT-B (85.8M) Qwen-Chat (7.1B) Q-Former (49.6M) 7.3B 16 50.5
    + NoteLLM-2 ViT-B (85.8M) Qwen-Chat (7.1B) Q-Former (49.6M) 7.3B 16 46.9
MQwen-bigG ViT-bigG (1.8B) Qwen-Chat (7.1B) Q-Former (142.9M) 9.1B 16 36.8
    + NoteLLM-2 ViT-bigG (1.8B) Qwen-Chat (7.1B) Q-Former (142.9M) 9.2B 16 35.2
Appendix F More Information about Saliency Scores
F.1 Formal Definition of $S_{v}$, $S_{t}$ and $S_{o}$
$S_{v}$, the mean significance of information flow from the vision embeddings $E_{v}$ to the final compressed position $c$:
$S_{v}$ $=\frac{\sum_{(i,j)\in P_{v}}I_{l}(i,j)}{\left|P_{v}\right|},$ (9)
$P_{v}$ =$=\{(c,j):j\in E_{v}\}.$
$S_{t}$, the mean significance of information flow from the textual embeddings $E_{t}$ to the final compressed position $c$:
$S_{t}$ $=\frac{\sum_{(i,j)\in P_{t}}I_{l}(i,j)}{\left|P_{t}\right|},$ (10)
$P_{t}$ =$=\{(c,j):j\in E_{t},j<c\}.$
$S_{o}$, the mean significance of information flow amongst all words, excluding influences represented by $S_{v}$ and $S_{t}$:
$S_{o}$ $=\frac{\sum_{(i,j)\in P_{o}}I_{l}(i,j)}{\left|P_{o}\right|},$ (11)
$P_{o}$ =$=\{(i,j):j<i\}-P_{v}-P_{t}.$
F.2 Results of Qwen-VL-Chat
Figure 5: Relative sizes of $S_{v}$, $S_{t}$ and $S_{o}$ in different layers of Qwen-VL-Chat.
We observe a unique pattern of saliency scores in Qwen-VL-Chat, as shown in Figure 5. The predominant information flow in Qwen-VL-Chat is textual, with a value higher than $0.8$ in all layers. This result is attributed to the extensive visual embeddings ($256$ in Qwen-VL-Chat) inputted into LLMs, which reduces the average saliency of image information flow. This denotes that traditional MLLMs are not directly suitable for MLRMs. The necessary action to adapt MLLMs to MLRMs is enhancing the efficiency of information flow.
Appendix G Performance of Zero-shot Multimodal Representation
Table 10: Zero-shot representation performance of different models (%).
  Method Input All Pair Short Query Pair Short Target Pair
R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
  BLIP-2 Image 19.60 34.16 53.49 19.08 34.92 56.18 20.06 36.41 58.20
Text 15.83 27.29 44.50 9.42 16.36 31.29 12.16 21.15 39.94
Multimodal 27.28 43.00 62.69 22.46 36.93 59.54 25.16 40.54 60.92
  LLaVA-1.5 Image 18.88 33.17 53.14 19.64 34.18 56.18 21.00 35.38 57.10
Text 30.86 46.71 64.14 21.90 34.43 49.63 22.28 32.98 47.48
Multimodal 37.01 55.46 73.71 30.51 47.78 68.73 31.56 48.09 68.13
  Qwen-VL Image 17.55 31.62 51.58 17.77 33.55 55.56 18.41 33.88 55.97
Text 38.74 54.65 70.37 23.31 35.17 51.95 22.17 34.84 51.29
Multimodal 37.88 55.51 74.73 26.28 43.20 64.89 29.82 47.63 69.72
  Qwen-VL-Chat Image 17.43 31.15 50.48 17.59 32.29 55.04 18.75 33.40 55.09
Text 37.92 53.02 68.48 22.63 33.87 49.15 21.32 32.50 48.03
Multimodal 40.93 58.59 75.44 29.10 44.48 64.14 31.81 49.09 68.57
  Qwen-Chat Text 38.98 54.08 69.02 23.08 34.90 50.44 26.04 38.35 53.33
Tomato Text 42.67 60.54 78.01 24.51 37.35 56.45 32.27 46.85 66.75
BM25 Text 55.20 72.67 82.83 35.82 48.42 59.71 35.07 49.63 63.68
In this section, we explore the multimodal representation ability of MLLMs in a zero-shot way in the multimodal I2I recommendation task. We select four popular MLLMs, i.e., BLIP-2 [25], LLaVA-1.5 [29], Qwen-VL [3] and Qwen-VL-Chat [3]. We also compare some baselines BM25 [40], Qwen-Chat [2] and our continual pre-trained LLM Tomato, which is pre-trained on our platform data based on LLaMA 2 [47], but lacks vision perception ability. The MLLMs details are concluded in Table 9. To understand the representation ability of MLLMs for different modalities, we also input images and texts independently for the representation test.
The results are presented in Table 10. We have several observations. Firstly, the zero-shot representation performance of existing MLLMs is inferior to BM25. This indicates that despite their excellent visual understanding, the zero-shot multimodal representation ability of MLLMs is unsatisfactory. This is because MLLMs are trained using language modeling loss, which mismatches representation tasks. Therefore, additional training is needed to properly align MLLMs with representation tasks. Next, we find that MLLMs can achieve better performance for multimodal input in most cases, which demonstrates that MLLMs can extract and fuse multimodal information effectively. At the same time, we discover that most MLLMs are better at representing text information than image information. This is reasonable because most parameters of MLLMs originate from LLMs, and the text of notes is more discriminative than images in our scenarios. Lastly, we find that LLMs can achieve performance comparable to MLLMs without any image input.
Appendix H Case Study
Figure 6: The visualization cases of some methods. Each line is a case, which the left two columns are the query note and target note, the other columns are the top-$1$ ranked notes for each method. We also show the target note ranking.
We show some cases for multimodal I2I note recommendation as shown in Figure 6. In the first case, both the query note and the target note discuss a guide to Oahu, an island in Hawaii. MTomato-Base + NoteLLM-2 successfully ranks the target note in the top position. While METER Co-attn and MTomato-Base also recall relevant notes in the first place, they differ slightly. One talks about an unpleasant experience in Oahu, and the other discusses an overall trip plan for Hawaii. The second example illustrates a note pair about the Pagani car model. Although all baselines recall notes about various car models, these notes discuss many different types of cars. MTomato-Base + NoteLLM-2 ranks the Pagani model car in the first place, which can be attributed to both the image and the text content. The third case discusses how to make good coffee and coffee tricks. LLM-based methods recall relevant notes about the tools needed to make coffee and how to do it. In contrast, METER Co-attn, which has a weaker understanding of text, only recommends coffee beans. The last case involves a short note pair. The query note is about an AI drawing for China’s National Day, and the target note is about a mooncake drawn by AI, which relates to the Chinese traditional festival, the Mid-Autumn Festival. We find that MTomato-Base + NoteLLM-2 can understand the query note image about Chinese culture and recommends notes talking about Chang’e, a figure from Chinese folklore associated with the Mid-Autumn Festival. Tomato and MTomato-Base only focus on the AI painting, rather than the image content.

### v2

@@
 NoteLLM-2: Multimodal Large Representation Models for Recommendation
-Abstract
-Large Language Models (LLMs) have demonstrated exceptional text understanding.
-Existing works explore their application in text embedding tasks.
-However, there are few works utilizing LLMs to assist multimodal representation tasks.
-In this work, we investigate the potential of LLMs to enhance multimodal representation in multimodal item-to-item (I2I) recommendations.
-One feasible method is the transfer of Multimodal Large Language Models (MLLMs) for representation tasks.
-However, pre-training MLLMs usually requires collecting high-quality, web-scale multimodal data, resulting in complex training procedures and high costs.
-This leads the community to rely heavily on open-source MLLMs, hindering customized training for representation scenarios.
-Therefore, we aim to design an end-to-end training method that customizes the integration of any existing LLMs and vision encoders to construct efficient multimodal representation models.
-Preliminary experiments show that fine-tuned LLMs in this end-to-end method tend to overlook image content.
-To overcome this challenge, we propose a novel training framework, NoteLLM-2, specifically designed for multimodal representation.
-We propose two ways to enhance the focus on visual information.
-The first method is based on the prompt viewpoint, which separates multimodal content into visual content and textual content.
-NoteLLM-2 adopts the multimodal In-Content Learning method to teach LLMs to focus on both modalities and aggregate key information.
-The second method is from the model architecture, utilizing a late fusion mechanism to directly fuse visual information into textual information.
-Extensive experiments have been conducted to validate the effectiveness of our method.
-1 Introduction
-Figure 1: Two strategies for using LLMs to enhance multimodal representations.
-With the development of the internet, most online platforms are full of multimodal information to drive user engagement [51, 50, 53].
-Multimodal recommendation is a necessary service for these platforms [30], which recommends items based on users’ interest on item multimodal information.
-These techniques mainly rely on multimodal representations, extracting raw multimodal content into dense embeddings [61], and then measuring similarity through these multimodal embeddings.
-The enhancement of multimodal representations can fundamentally improve numerous downstream tasks, such as retrieval [31], clustering [34], and item modeling [60, 51].
-Many existing works have investigated multimodal representations, such as CLIP [37], METER [14] and Siglip [58].
-Although great success has been achieved in these domains, the number of parameters in most multimodal representation models still possesses the large potential for scaling [45, 46].
-However, existing works [45, 46] primarily focus on scaling vision models to improve multimodal representation, overlooking the textual branch.
-At the same time, due to the lack of sufficient pure text-oriented pre-training [37, 45, 46, 58], the text branch in multimodal representation models often exhibits suboptimal understanding of text information [6, 59].
-Therefore, it is necessary to enhance text understanding in multimodal representations models.
-Recently, Large Language Models (LLMs) have shown impressive capabilities in text understanding and generating [18, 7, 2, 47].
-Several works show the superior performance of LLMs in generating textual embeddings [34, 60, 19, 31, 4].
+Abstract.
+Large Language Models (LLMs) have demonstrated exceptional proficiency in text understanding and embedding tasks.
+However, their potential in multimodal representation, particularly for item-to-item (I2I) recommendations, remains underexplored.
+While leveraging existing Multimodal Large Language Models (MLLMs) for such tasks is promising, challenges arise due to their delayed release compared to corresponding LLMs and the inefficiency in representation tasks.
+To address these issues, we propose an end-to-end fine-tuning method that customizes the integration of any existing LLMs and vision encoders for efficient multimodal representation.
+Preliminary experiments revealed that fine-tuned LLMs often neglect image content.
+To counteract this, we propose NoteLLM-2, a novel framework that enhances visual information.
+Specifically, we propose two approaches: first, a prompt-based method that segregates visual and textual content, employing a multimodal In-Context Learning strategy to balance focus across modalities; second, a late fusion technique that directly integrates visual information into the final representations.
+Extensive experiments, both online and offline, demonstrate the effectiveness of our approach.
+Code is available at https://github.com/Applied-Machine-Learning-Lab/NoteLLM.
+Multimodal Large Language Model; Recommendation; Multimodal Representation
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709440isbn: 979-8-4007-1245-6/25/08ccs: Information systems Recommender systems
+1. Introduction
+Figure 1.
+Two strategies for using LLMs to enhance multimodal representations.
+The first strategy transfers existing pre-trained MLLMs into representation tasks.
+The second strategy fine-tunes the integration of LLMs and vision encoders end-to-end, requiring no alignment and offering better practicality and efficiency.
+With the development of the internet, most online platforms are full of multimodal information to drive user engagement (Wu et al., 2022; Wei et al., 2019; Xun et al., 2021).
+Multimodal recommendation is a necessary service for these platforms (Liu et al., 2023a), which recommends items based on users’ interest in item multimodal information.
+These techniques mainly rely on multimodal representations, extracting raw multimodal content into dense embeddings (Zhang et al., 2020), and then measuring similarity through these multimodal embeddings.
+The enhancement of multimodal representations can fundamentally improve numerous downstream tasks, such as retrieval (Ma et al., 2023), clustering (Muennighoff et al., 2024), and item modeling (Wu et al., 2022).
+Many existing efforts have investigated multimodal representations, such as CLIP (Radford et al., 2021) and METER (Dou et al., 2022).
+Despite significant success in these areas, the number of parameters in most multimodal representation models still has great potential for scaling (Sun et al., 2023, 2024).
+However, existing works (Sun et al., 2023, 2024) primarily focus on scaling vision models to improve multimodal representation, overlooking the textual branch.
+At the same time, due to the lack of sufficient pure text-oriented pre-training (Radford et al., 2021; Sun et al., 2023, 2024; Zhai et al., 2023), the text branch in multimodal representation models often exhibits suboptimal understanding of text information (Chen et al., 2023; Zhang et al., 2024c).
+Therefore, it is necessary to enhance text understanding in multimodal representation models.
+Recently, Large Language Models (LLMs) have shown impressive capabilities in text understanding and generating (Jiang et al., 2023b; Chiang et al., 2023; Bai et al., 2023a; Touvron et al., 2023).
+Several works show the superior performance of LLMs in generating textual embeddings (Muennighoff et al., 2024; Zhang et al., 2024a; Jiang et al., 2023a; Ma et al., 2023; BehnamGhader et al., 2024).
 However, the use of LLMs for enhancing text understanding in multimodal representations is underexplored.
-As shown in Figure 1, one feasible method is to adapt Multimodal Large Language Models (MLLMs) [29, 25, 2, 22] for representation tasks.
-Nevertheless, pre-training MLLMs requires collecting high-quality, web-scale multimodal data, resulting in complex training procedures and high costs [25, 2].
-This leads the community to rely heavily on open-source MLLMs, hindering customized training using the most advanced LLMs and vision encoders for representation scenarios.
-Therefore, we aim to design an end-to-end training method that customizes the integration of any existing LLMs and vision encoders to construct efficient multimodal representation models.
-This end-to-end method is practical when the target LLMs lack pre-trained visual components [11, 18, 1], as it only needs target domain data, eliminating pre-training data collection.
+As shown in Figure 1, one feasible method is to adapt existing open-source Multimodal Large Language Models (MLLMs) (Liu et al., 2023b; Li et al., 2023; Bai et al., 2023a; Karamcheti et al., 2024) for representation tasks.
+However, developing MLLMs requires substantial high-quality multimodal data and significant computational resources for pre-training (Li et al., 2023; Bai et al., 2023a), resulting in the release of LLMs without multimodal capabilities ahead of their multimodal counterparts (DeepSeek-AI, 2024; Jiang et al., 2023b; lla, [n.
+d.]).
+This delay constrains customized training using the most advanced LLMs and vision encoders.
+Moreover, existing MLLMs are inefficient for representation scenarios as they primarily focus on multimodal understanding and generation (Shang et al., 2024).
+Therefore, we design an end-to-end fine-tuning method that customizes the integration of any existing LLMs and vision encoders to construct efficient multimodal representation models.
+This end-to-end method is practical when the target LLMs lack pre-trained visual components (DeepSeek-AI, 2024; Jiang et al., 2023b; lla, [n.
+d.]), as it only needs target domain data, eliminating pre-training data collection.
 Besides, we can optimize the customized model structure for representation tasks to achieve better efficiency while ensuring performance.
-We call these models for enhancing multimodal representation with LLMs as Multimodal Large Representation Models (MLRMs).
+We refer to these models for enhancing multimodal representation with LLMs as Multimodal Large Representation Models (MLRMs).
+Figure 2.
+The frameworks of MLRMs.
+(a) is the basic method for representations.
+(b) is our representation method, which contains two easy and effective mechanisms to enhance multimodal representation ability, mICL and late fusion.
 In this work, we explore MLRMs in item-to-item (I2I) multimodal recommendation scenarios.
-We conduct fine-tuned preliminary experiments to examine our goal.
-Our findings reveal that representations of naive fine-tuned end-to-end MLRMs are biased towards textual information, ignoring visual content.
-Visual information is mainly used to amplify the key element in multimodal content, rather than being directly aggregated by representations.
-To solve this challenge, we propose a novel training framework, called NoteLLM-2, for multimodal representation tasks.
-We design two methods to mitigate visual modality neglect.
+We conduct preliminary fine-tuning experiments to examine our goal.
+Our findings reveal that naive end-to-end fine-tuned MLRMs are biased towards text, ignoring visual content.
+Images are mainly used to amplify key elements in multimodal content, rather than being directly aggregated by representations.
+To solve this challenge, we propose NoteLLM-2, a novel fine-tuning framework for multimodal tasks, with two methods to mitigate visual neglect.
 The first method is multimodal In-Context Learning (mICL).
@@
 Each modality-compressed word conducts in-batch contrastive learning with its corresponding modality-compressed word.
-During inference, mICL can aggregate multimodal information in an In-Context Learning (ICL) way [12, 48].
 The second method is the late fusion mechanism.
-Most MLLMs adopt early fusion, which maps visual features into the textual feature space of LLMs and interact visual features with text embeddings through LLMs [29, 25, 2, 22].
-Even though early fusion can deepen the multimodal information interaction, it can easily lead to the loss of pure visual information during the multimodal fusion conducted by LLMs [25, 62].
+Most MLLMs adopt early fusion, which maps visual features into the textual feature space of LLMs and interacts visual features with text embeddings through LLMs (Liu et al., 2023b; Li et al., 2023; Bai et al., 2023a; Karamcheti et al., 2024).
+Even though early fusion can deepen the multimodal information interaction, it can easily lead to the loss of pure visual information during the multimodal fusion conducted by LLMs (Li et al., 2023; Zhang et al., 2024b).
 We incorporate late fusion based on early fusion, retaining more visual information by delaying the fusion process.
-Specifically, we adopt a multimodal gate fusion mechanism [8], which takes visual representations from the visual encoder and representations from the LLM to generate final multimodal representations.
+Specifically, we adopt a multimodal gate fusion mechanism, which takes visual representations from the visual encoder and representations from the LLM to generate final multimodal representations.
 We conduct extensive experiments to demonstrate our NoteLLM-2 effectiveness.
 Our paper makes the following contributions:
-  • To the best of our knowledge, we are the first to explore the multimodal representation assisted by LLMs.
-We propose an end-to-end training method that customizes the integration of any existing LLMs and vision encoders to form efficient MLRMs.
-  • We reveal that fine-tuning unaligned MLRMs for representation tasks can cause the modality imbalance, overlooking visual information.
-To overcome this problem, we design a novel framework, NoteLLM-2, which contains two mechanisms, mICL and late fusion.
-  • Extensive experiments have demonstrated the effectiveness of our proposed framework for multimodal recommendation representation.
-Our paper is organized as follows: We begin with a preliminary investigation of MLRMs in multimodal I2I recommendation scenarios and analyze the results.
-Then, inspired by this analysis, we design our NoteLLM-2.
-And we conduct experiments to demonstrate the effectiveness of our NoteLLM-2.
-Besides, related works and additional details are presented in Appendix A.
-2 Preliminary Investigation
-2.1 Preliminaries
-Problem Statement.
+  • To the best of our knowledge, we are the first to explore the multimodal representation assisted by LLMs in recommendation scenarios.
+We propose an end-to-end fine-tuning method that customizes the integration of any existing LLMs and vision encoders to form efficient MLRMs.
+  • We reveal that fine-tuning unaligned MLRMs for representation tasks can cause modality imbalance by overlooking visual information.
+To overcome this problem, we design a novel framework, NoteLLM-2, which contains mICL and late fusion.
+  • Extensive online and offline experiments have demonstrated the effectiveness of our proposed framework for multimodal recommendation representation.
+2. Preliminary Investigation
+2.1. Problem Statement
 In our scenarios, every item represents a note, which is generated by users and expresses their life experiences.
-We show some notes in Appendix H.
 We assume $\mathcal{N}=\{n_{1},n_{2},...,n_{m}\}$ as note pool, where $m$ is the number of notes.
@@
 The objective of this system is to prioritize the ranking of the related target note.
-Dataset Construction.
-Following [60], we utilize the co-occurrence mechanism to build related note pairs based on user behaviors without human annotations.
-This mechanism assumes that notes frequently read together are likely related.
-Then, we count the co-occurrence score in which users viewed note $n_{A}$ and subsequently clicked on note $n_{B}$ as follows:
-$s_{n_{A}\rightarrow n_{B}}=\sum_{u\in U_{n_{A}\rightarrow n_{B}}}\frac{1}{N_{u}},$ (1)
-where $s_{n_{A}\rightarrow n_{B}}$ means the co-occurrence score from note $n_{A}$ to note $n_{B}$, $U_{n_{A}\rightarrow n_{B}}$ is the set of users who viewed note $n_{A}$ and subsequently clicked on note $n_{B}$, and $N_{u}$ denotes the quantity of the note set clicked by the user $u$ in the user behavior data.
-After calculating the co-occurrence score for all note pairs, we construct the set of co-occurrence scores $\mathcal{S}_{n_{i}}$ from note $n_{i}$ to all other notes.
-Specifically, $\mathcal{S}_{n_{i}}$ is defined as $\{{s_{n_{i}\rightarrow n_{j}}}|1\leq j\leq m,i\neq j\}$.
-Next, we remove outlier from $\mathcal{S}_{n_{i}}$ with co-occurrence scores above $up$ or below $low$.
-Finally, we select the $t$ notes with the highest co-occurrence scores from the filtered set as the related notes for note $n_{i}$.
-Figure 2: The frameworks of MLRMs.
-(a) is the basic method for representations.
-(b) is our representation method, which contains two easy and effective mechanisms to enhance multimodal representation ability, mICL and late fusion.
-Basic Representation Method of MLRMs.
-Following [19], we adopt a prompt with explicit one word limitation to compress the multimodal content into one embedding.
+2.2. Basic Representation Method of MLRMs
+Following (Jiang et al., 2023a), we adopt a prompt with an explicit one-word limitation to compress the multimodal content into one embedding.
 Specifically, we use json format to show note content and our prompt is the following: {mdframed} Note Compression Prompt
-Note content: {’image’: <IMG>, ’title’: $t_{i}$ , ’topic’: $tp_{i}$ , ’content’: $ct_{i}$}.
-Compress this note into one word:" In this template, <IMG> is a placeholder, which will be replaced by vision embeddings processed from raw images $v_{i}$.
+Note content: {‘image’: ¡IMG¿, ‘title’: $t_{i}$ , ‘topic’: $tp_{i}$ , ‘content’: $ct_{i}$}.
+Compress this note into one word:“ In this template, ¡IMG¿ is a placeholder, which will be replaced by vision embeddings processed from raw images $v_{i}$.
 The overall framework is shown in Figure 2 (a).
@@
 After processing images, the text prompts are tokenized into discrete indexes, then form the text word embedding $E_{t}\in\mathbb{R}^{T\times h_{t}}$, where $T$ is the length of text tokens.
-To insert the visual embeddings into the correct position in text embeddings, we replace word embeddings in positions of the <IMG> token by the visual embeddings $E_{v}$, forming the multimodal embeddings $E_{m}\in\mathbb{R}^{(L_{c}+T-1)\times h_{t}}$.
+To insert the visual embeddings into the correct position in text embeddings, we replace word embeddings in positions of the ¡IMG¿ token by the visual embeddings $E_{v}$, forming the multimodal embeddings $E_{m}\in\mathbb{R}^{(L_{c}+T-1)\times h_{t}}$.
 Last, we use LLMs $LLM_{\mu}$ to process the multimodal embeddings to generate last hidden states $H\in\mathbb{R}^{(L_{c}+T-1)\times h_{t}}$, where $H=LLM_{\mu}(E_{m})$.
-We take last embeddings of $H$ as the note representation $\boldsymbol{n}_{i}$ for note $n_{i}$.
+We take the last embeddings of $H$ as the note representation $\boldsymbol{n}_{i}$ for note $n_{i}$.
 This method constrains LLMs to compress the multimodal content into one embedding in the form of the next token prediction.
-However, LLMs are trained using the language modeling loss [29, 25, 3, 22], which significantly differs from representation tasks.
-To bridge this gap, we employ contrastive learning, a prevalent technique for training embeddings [34, 60, 19, 31].
-Specifically, each minibatch contains $B$ related note pairs $(n_{i},n_{i}^{+})$, resulting in a total of $2B$ notes per minibatch.
-Following [36], we use gradient descent to minimize the contrastive loss as follows:
-$\mathcal{L}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\boldsymbol{n}_{i},\boldsymbol{n}_{i}^{+})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\boldsymbol{n}_{i},\boldsymbol{n}_{j})\cdot e^{\tau}}},$ (2)
+However, LLMs are trained using the language modeling loss (Liu et al., 2023b; Li et al., 2023; Bai et al., 2023b; Karamcheti et al., 2024), which significantly differs from representation tasks.
+To bridge this gap, we employ contrastive learning, a prevalent technique for training embeddings (Muennighoff et al., 2024; Zhang et al., 2024a; Jiang et al., 2023a; Ma et al., 2023).
+Specifically, each minibatch contains $B$ related note pairs, resulting in a total of $2B$ notes per minibatch.
+For any note $n_{i}$ (where $1\leq i\leq 2B$) in the minibatch, we denote its related note in the same minibatch as $n_{i}^{+}$.
+Following (Neelakantan et al., 2022), we use gradient descent to minimize the contrastive loss as follows:
+(1) $\mathcal{L}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\boldsymbol{n}_{i},\boldsymbol{n}_{i}^{+})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\boldsymbol{n}_{i},\boldsymbol{n}_{j})\cdot e^{\tau}}},$
 where $\tau$ means the learnable temperature and $sim(a,b)=a^{\top}b/(\|a\|\|b\|)$.
-$\mathcal{L}(\pi,\mu)$ means that we only update the connector $C_{\pi}$ and LLMs $LLM_{\mu}$, while keep the vision encoder $V_{\theta}$ frozen, aiming for the larger batch size and better performance [22].
-2.2 Datasets and Experimental Settings
-Datasets.
-We obtain a real-world multimodal I2I dataset from our platform.
-We create the training dataset by randomly choosing related note pairs from user behaviour data gathered over two weeks.
-One-tenth of pairs in this training dataset is used for the validation set.
-Next, we randomly select notes from the upcoming week to form the note pool of testing set, excluding any notes that are already in the training dataset.
-Then, we count the related note pairs in the note pool of testing set.
-The training dataset contains $1.5M$ notes and $1.1M$ note pairs.
-The testing dataset contains $0.5M$ notes and $21K$ note pairs.
-We provide detailed statistics of the training and testing dataset in Appendix C.
-To better evaluate the multimodal representation ability of MLRMs rather than relying heavily on the text modality, we collect pairs that include short notes in the test dataset.
-We define notes with a token length of less than $50$ as short notes, which approximately constitute $10\%$ of the total test notes.
-We categorize pairs with short query notes as short query pairs and those with short target notes as short target pairs.
-In the test dataset, the number of short query pairs is $5,620$ and the number of short target pairs is $5,582$.
-Experimental settings.
-In related note pair construction, we set the upper bound of the co-occurrence score $up$ as $30$ and the lower bound $low$ as $0.01$.
-And we set $t$ as $3$.
-To maintain context length restrictions, we truncate the titles exceeding $20$ tokens, and truncate the contents exceeding $80$ tokens.
-In the fine-tuned experiments, for fair comparison, we add a linear projector to diminish the dimension of note embeddings to $64$.
-The batch size $B$ is set to 128128128128111We train models on 8 $\times$ 80GB Nvidia A100 GPUs and the batch size per each GPU is $16$..
-Each batch contains $256$ notes.
-The temperature $\tau$ is initialized as $3$.
-We provide more training details in Appendix D.
-To evaluate the representation ability of MLRMs, we rank all notes in the note pool, excluding the query note, according to the content of query note.
-We compute recall scores based on the position of the target note in the ranked list.
-We report Recall@100, Recall@1k and Recall@10k on all pairs, short query pairs and short target pairs in the test dataset.
-Table 1: Representation performance of different models (%).
+$\mathcal{L}(\pi,\mu)$ means that we only update the connector $C_{\pi}$ and LLMs $LLM_{\mu}$, while keep the vision encoder $V_{\theta}$ frozen, aiming for the larger batch size and better performance (Karamcheti et al., 2024).
+Table 1. Detailed statistics of training and testing dataset.
+                                  training dataset
+  # notes 1,487,613 # note pairs 1,130,906
+avg. # words per title 12.38 avg. # topic per note 5.26
+avg. # words per topic 4.53 avg. # words per content 133.20
+                                  testing dataset
+  # notes 534,144 # note pairs 21,262
+avg. # words per title 11.68 avg. # topic per note 4.58
+avg. # words per topic 4.50 avg. # words per content 116.00
+Table 2. Performance of different fine-tuned models (%). “*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline.
   Method Input All Pair Short Query Pair Short Target Pair
@@
   CLIP ViT-B Image 23.21 45.44 71.16 25.24 49.40 75.99 25.58 50.28 75.80
-BM25 Text 55.20 72.67 82.83 35.82 48.42 59.71 35.07 49.63 63.68
+BM25 Text 55.22 72.64 82.82 35.75 48.41 59.81 35.05 49.52 63.64
 RoBERTa-wwm-ext Text 67.59 85.06 93.86 41.23 58.61 76.40 41.80 60.65 80.04
@@
 MQwen-Base Multimodal 74.02 89.65 97.22 48.15 68.35 88.22 49.82 70.91 91.43
-MQwen-bigG Multimodal 77.64 92.89 98.91 57.45 80.92 95.73 57.90 83.49 96.99
-Qwen-VL-Chat [^2] Due to Qwen-VL-Chat having $256$ visual embeddings per image, this model is trained on 32 $\times$ 80GB Nvidia A100 GPUs and the batch size per each GPU is $4$.
-Multimodal 78.54 93.77 99.03 60.43 83.26 96.60 61.63 85.48 97.89
-2.3 Performance of Multimodal Representation of Fine-tuned MLRMs
-We conduct zero-shot experiments on several existing MLLMs in Appendix G, and find that zero-shot is insufficient to adapt MLLMs to representation tasks, performing inferiorly compared to a simple baseline, BM25 [40].
-This is due to the misalignment between pre-training tasks and representation tasks [60, 19].
+MQwen-bigG Multimodal 77.63 92.89 98.91 57.45 80.92 95.73 57.95 83.50 96.99
+Qwen-VL-Chat [^1] Due to Qwen-VL-Chat having $256$ visual embeddings per image, this model is trained on 32 $\times$ 80GB Nvidia A100 GPUs and the batch size per GPU is $4$.
+Multimodal 78.53* 93.76* 99.03* 60.41* 83.27* 96.60* 61.60* 85.48* 97.89*
+2.3. Datasets and Experimental Settings
+Dataset Construction Method.
+To construct pairs of related notes without relying on human annotations, we leverage the co-occurrence mechanism (Zhang et al., 2024a), which measures the relevance between notes by how often they are read together.
+We count the co-occurrence score in which users viewed note $n_{A}$ and subsequently clicked on note $n_{B}$ as follows:
+(2) $s_{n_{A}\rightarrow n_{B}}=\sum_{u\in U_{n_{A}\rightarrow n_{B}}}\frac{1}{N_{u}},$
+where $s_{n_{A}\rightarrow n_{B}}$ means the co-occurrence score from note $n_{A}$ to note $n_{B}$, $U_{n_{A}\rightarrow n_{B}}$ is the set of users who viewed note $n_{A}$ and subsequently clicked on note $n_{B}$, and $N_{u}$ denotes the number of distinct notes clicked by the user $u$ in the user behavior data.
+We compute the co-occurrence score for each note pair.
+Then, we construct the set $\mathcal{S}_{n_{i}}$, which contains co-occurrence scores from note $n_{i}$ to all other notes, i.e., $\mathcal{S}_{n_{i}}=\{{s_{n_{i}\rightarrow n_{j}}}|1\leq j\leq m,i\neq j\}$.
+Next, we remove outlier from $\mathcal{S}_{n_{i}}$ with co-occurrence scores above $up$ or below $low$.
+Finally, the top $t$ notes with the highest co-occurrence scores in the filtered set are viewed as the related notes for note $n_{i}$.
+Datasets Details.
+We obtain a real-world multimodal I2I dataset from our platform.
+We create the training dataset by randomly choosing related note pairs from user behavior data gathered over two weeks.
+One-tenth of the pairs in this training dataset are used for the validation set.
+Next, the test note pool is collected by randomly selecting notes from the following week, while avoiding duplication with the training dataset.
+Then, we count the related note pairs in the test note pool.
+We provide detailed statistics of the training and testing dataset in Table 1.
+To better evaluate the multimodal representation ability of MLRMs rather than relying heavily on the text, we collect pairs that include short notes in the test dataset.
+We define notes with a token length of less than $50$ as short notes, which approximately constitute $10\%$ of the total test notes.
+We categorize pairs with short query notes as short query pairs and those with short target notes as short target pairs.
+In the test dataset, the number of short query pairs is $5,620$ and the number of short target pairs is $5,582$.
+Experimental settings.
+When constructing pairs of related notes, the upper bound of the co-occurrence score $up$ is $30$ and the lower bound $low$ is $0.01$.
+We set $t$ as $3$.
+We truncate titles longer than 20 words and content longer than 80 words to comply with inference length restrictions.
+In the fine-tuned experiments, for fair comparison, we add a linear projector to diminish the dimension of note embeddings to $64$.
+The batch size $B$ is set to 128128128128111Models trained on 8 $\times$ 80GB Nvidia A100 GPUs and the batch size per GPU is $16$..
+There are $256$ notes in each batch.
+We start with a temperature $\tau$ set as $3$.
+We provide more training details in Appendix A.
+To evaluate the representation ability for recall tasks, we create a similarity ranked list of all remaining notes in the test pool based on the query note’s content.
+We compute recall scores based on the position of the target note in the ranked list.
+We report Recall@100, Recall@1k, and Recall@10k on all pairs, short query pairs, and short target pairs in the test dataset.
+For robustness, we report the average results tested three times using random seeds {42,43,44}.
+2.4. Performance of Multimodal Representation of Fine-tuned MLRMs
+We conduct zero-shot experiments on several existing MLLMs in Appendix D, and find that zero-shot is insufficient to adapt MLLMs to representation tasks, performing inferiorly compared to a simple baseline, BM25 (Robertson et al., 2009).
+This is due to the misalignment between pre-training tasks and representation tasks (Zhang et al., 2024a; Jiang et al., 2023a).
 Hence, fine-tuning MLLMs for representation tasks is necessary.
-We design three end-to-end MLRMs to examine the representation training method: MTomato-Base uses Tomato (our continually pre-trained LLM based on LLaMA 2 [47], which lacks vision perception ability) as the LLM, CLIP ViT-B [37] as the vision encoder and a randomly initialized Q-Former [25] as the connector for efficiency.
-MQwen-Base replaces Tomato with Qwen-Chat [2] in MTomato-Base.
-MQwen-bigG replaces CLIP ViT-B with ViT-bigG [17] in MQwen-Base.
-We set the visual embedding length of these models as $16$ for efficiency.
-For comparison, we choose two pre-trained MLLMs: BLIP-2 and Qwen-VL-Chat [3].
-All vision encoders in these models are frozen for the larger batch size.
-In addition to MLRMs, we also compare other existing multimodal representation baselines.
-We provide details about baselines and MLRMs in Appendix E.
-The results are shown in Table 1.
-Several interesting observations can be made.
-Firstly, MLRMs based on LLMs can outperform existing baselines significantly.
-Qwen-VL-Chat achieves a $10.78\%$ improvement in R@100 compared to the traditional method, METER Co-attn on all pairs.
-Secondly, the end-to-end training representation method can enhance the model’s multimodal representation ability.
-MQwen-bigG improves R@100 on all pairs by $6.31\%$ compared to Qwen-Chat.
-However, when the visual encoder is small, such as CLIP ViT-B, the enhancement in multimodal perception is not substantial.
-MTomato-Base only improves R@100 on all pairs by $1.31\%$ compared to Tomato, and MQwen-Base only improves R@100 on all pairs by $1.36\%$ compared to Qwen-Chat.
-Lastly, although MQwen-bigG is more efficient than Qwen-VL-Chat, which uses the same vision encoder and LLM (see Appendix E.3), there remains a performance gap.
-Table 2: Representation performance of fine-tuned MLRMs under different modal inputs (%).
+Table 3. Performance of MLRMs fine-tuned with multimodal inputs under different modal inputs (%).
   Method Input All Pair Short Query Pair Short Target Pair
@@
 Text 70.63 86.09 93.81 39.94 55.77 72.17 43.03 60.02 79.77
-Multimodal 77.64 92.89 98.91 57.45 80.92 95.73 57.90 83.49 96.99
+Multimodal 77.63 92.89 98.91 57.45 80.92 95.73 57.95 83.50 96.99
   Qwen-VL-Chat Image 24.67 45.51 70.08 25.94 48.32 75.32 26.94 49.59 75.67
 Text 71.44 86.78 94.43 42.55 58.59 75.86 43.97 61.11 80.17
-Multimodal 78.54 93.77 99.03 60.43 83.26 96.60 61.63 85.48 97.89
-We also evaluate all fine-tuned MLRMs under different modal inputs, as shown in Table 2.
-Firstly, we observe that MTomato-Base and MQwen-Base hardly represent the image content when the input consists solely of images, which shows ignoring visual information.
+Multimodal 78.53 93.76 99.03 60.41 83.27 96.60 61.60 85.48 97.89
+We design three end-to-end MLRMs to examine the representation fine-tuning method: MTomato-Base uses Tomato (our continually pre-trained LLM based on LLaMA 2 (Touvron et al., 2023), which lacks vision perception ability.
+See Appendix B for more details on pre-training.) as the LLM, CLIP ViT-B (Radford et al., 2021) as the vision encoder and a randomly initialized Q-Former (Li et al., 2023) as the connector for efficiency.
+MQwen-Base replaces Tomato with Qwen-Chat (Bai et al., 2023a) in MTomato-Base.
+MQwen-bigG replaces CLIP ViT-B with ViT-bigG (Ilharco et al., 2021) in MQwen-Base.
+We set the visual embedding length of these models as $16$ for efficiency.
+For comparison, we choose two pre-trained MLLMs: BLIP-2 and Qwen-VL-Chat (Bai et al., 2023b).
+All vision encoders in these models are frozen for the larger batch size.
+We provide details about end-to-end MLRMs in Appendix B.
+To evaluate the effectiveness of MLRMs, we compare them with several baselines.
+(1) CLIP ViT-B (Radford et al., 2021) is trained using only image input.
+(2) BM25 (Robertson et al., 2009) is a basic baseline, which counts the relevance between the query and documents based on every word in the query.
+(3) RoBERTa-wwm-ext (Cui et al., 2021) is a classic traditional Chinese BERT-like (Kenton and Toutanova, 2019) text encoder with pure text input.
+(4) Tomato is our continually pre-trained LLM that is based on LLaMA 2 (Touvron et al., 2023).
+(5) NoteLLM (Zhang et al., 2024a) unifies the embedding generation and topic generation into one task.
+(6) METER Merge-attn and METER Co-attn (Dou et al., 2022) leverage best training strategy to train vision-and-language transformers.
+METER Merge-attn adopts the merged attention as the fusion module, while METER Co-attn uses co-attention as the fusion module.
+All baselines are trained on our training dataset.
+The results are shown in Table 2.
+Several observations can be made.
+Firstly, MLRMs can significantly outperform existing non-LLM-based baselines.
+Qwen-VL-Chat achieves a $10.78\%$ improvement in R@100 compared to the traditional method, METER Co-attn on all pairs.
+Secondly, the end-to-end fine-tuned representation method can enhance the model’s multimodal representation ability.
+MQwen-bigG improves R@100 on all pairs by $6.31\%$ compared to Qwen-Chat.
+However, when the visual encoder is small, such as CLIP ViT-B, the enhancement in multimodal perception is not substantial.
+MTomato-Base only improves R@100 on all pairs by $1.31\%$ compared to Tomato, and MQwen-Base only improves R@100 on all pairs by $1.36\%$ compared to Qwen-Chat.
+Lastly, although MQwen-bigG is more efficient than Qwen-VL-Chat, which uses the same vision encoder and LLM (see Section 4.4), there remains a performance gap.
+We also evaluate all MLRMs fine-tuned with multimodal inputs under different modal inputs, as shown in Table 3.
+Firstly, we observe that MTomato-Base and MQwen-Base hardly represent the image content when the input consists solely of images, which ignores visual information.
 Owing to its more robust vision encoder, MQwen-bigG exhibits superior vision representation capabilities.
 However, its performance still falls short when compared to BLIP-2 and Qwen-VL-Chat, both of which have undergone multimodal pre-training.
-2.4 Exploring the Bias of Fine-tuned MLRMs
-To further mine the bias in fine-tuned MLRMs, we explore the attention patterns of two modality tokens in different LLMs’ layers.
-Following [48], we adopt saliency scores to analyze the interaction between the compressed word and two modality tokens in the attention matrices.
+2.5. Exploring the Information Flow of Fine-tuned MLRMs
+To further explore the reasons why end-to-end fine-tuned MLRMs overlook visual information, we examine the attention patterns of two modality tokens in different LLMs’ layers from the perspective of information flow.
+We adopt saliency scores (Wang et al., 2023b) to analyze the interaction between the compressed word and two modality tokens in the attention matrices.
 We randomly select $10,000$ notes in batches from the training dataset to count the saliency scores, and compute the saliency scores as follows:
-$I_{l}=\sum_{h}\left|A^{\top}_{h,l}\frac{\partial\mathcal{L}}{\partial A_{h,l}}\right|,$ (3)
-where $A_{h,l}$ means the attention matrix value of the $h$-th attention head in the $l$-th layer and $\mathcal{L}$ is the contrastive loss.
-We sum all attention heads to obtain the saliency matrix $I_{l}$ of $l$-th layer.
+(3) $I_{l}=\sum_{h}\left|A^{\top}_{h,l}\frac{\partial\mathcal{L}}{\partial A_{h,l}}\right|,$
+where $A_{h,l}$ means the attention matrix value from the $h$-th attention head within $l$-th layer and $\mathcal{L}$ is the contrastive loss.
+The saliency matrix $I_{l}$ for layer $l$ is computed by summing across all attention heads.
 $I_{l}(i,j)$ means the significance of the information flow from the $j$-th word to the $i$-th word.
-To analyze different modality importances for the compressed word, we separate $I_{l}$ into three parts as flowing: $S_{v}$, the mean significance of information flow from the vision embeddings $E_{v}$ to the final compressed position $c$.
-$S_{t}$, the mean significance of information flow from the textual embeddings $E_{t}$ to the final compressed position $c$.
-$S_{o}$, the mean significance of information flow amongst all words, excluding influences represented by $S_{v}$ and $S_{t}$.
-We provide a more clear formal definition in Appendix F.1.
-The results of three parts from the saliency matrix are shown in Figure 3.
-MLRMs trained end-to-end without pre-training exhibit a low $S_{v}$ and a high $S_{o}$ in shallow layers, while pre-trained MLLMs, BLIP-2 and Qwen-VL-Chat (see Appendix F.2), have more balanced $S_{v}$ and $S_{o}$ in shallow layers.
-This indicates that end-to-end trained MLRMs primarily utilize the information gain from multimodal data to enhance the information flow in the context.
-Images in these models primarily assist the information flow in identifying crucial information [44, 43], rather than being directly aggregated by representations.
-Besides, text information flow significantly surpasses visual information flow in end-to-end trained MLRMs.
-((a)) BLIP-2
-((b)) MTomato-Base
-((c)) MQwen-Base
-((d)) MQwen-bigG
-Figure 3: Relative sizes of $S_{v}$, $S_{t}$ and $S_{o}$ in different layers of different MLRMs.
-3 Methodology
+To analyze the relative importance of different modalities for the compressed word, we separate $I_{l}$ into three parts as follows: $S_{v}$, the mean significance of information flow from the vision embeddings $E_{v}$ to the final compressed position $c$:
+(4) $S_{v}$ $=\frac{\sum_{(i,j)\in P_{v}}I_{l}(i,j)}{\left|P_{v}\right|},$
+$P_{v}$ =$=\{(c,j):j\in E_{v}\}.$
+$S_{t}$, the mean significance of information flow from the textual embeddings $E_{t}$ to the final compressed position $c$:
+(5) $S_{t}$ $=\frac{\sum_{(i,j)\in P_{t}}I_{l}(i,j)}{\left|P_{t}\right|},$
+$P_{t}$ =$=\{(c,j):j\in E_{t},j<c\}.$
+$S_{o}$, the mean significance of information flow amongst all words, excluding influences represented by $S_{v}$ and $S_{t}$:
+(6) $S_{o}$ $=\frac{\sum_{(i,j)\in P_{o}}I_{l}(i,j)}{\left|P_{o}\right|},$
+$P_{o}$ =$=\{(i,j):j<i\}-P_{v}-P_{t}.$
+The results of three parts from the saliency matrix are shown in Figure LABEL:fig:saliency_score.
+We observe that pre-trained MLLMs, BLIP-2 and Qwen-VL-Chat (see Appendix C), have balanced $S_{v}$ and $S_{o}$ in shallow layers.
+However, MLRMs fine-tuned end-to-end without pre-training exhibit a low $S_{v}$ and a high $S_{o}$ in shallow layers.
+The low $S_{v}$ indicates limited direct information flow from visual to final compressed embeddings, while the high $S_{o}$ indicates significant information flow between visual and textual embeddings.
+This shows that the information flow gain brought by introducing images is more likely to flow towards important multimodal embeddings, rather than being directly aggregated into final compressed embeddings.
+As the number of layers increases, we observe that $S_{v}$ increases while $S_{o}$ decreases dramatically.
+This is because, after multimodal embeddings capture context information in shallow layers (as shown by $S_{o}$), models focus on directly aggregating relevant information into final compressed embeddings for representation tasks.
+Unlike pre-trained MLLMs, the final compressed embeddings of end-to-end MLRMs have little direct visual information in shallow layers.
+This causes the final compressed embedding initially accumulates visual information mainly through the aggregation of context information contained in textual embeddings.
+In deep layers, final compressed embeddings increasingly incorporate image information related to the previously aggregated textual information through the attention mechanism.
+Therefore, images in end-to-end MLRMs primarily assist the information flow in identifying crucial information (Sun et al., 2021, 2020), rather than being directly aggregated by representations.
+Besides, text information flow significantly surpasses visual information flow in end-to-end fine-tuned MLRMs.
+Table 4. Performance of different methods based on three MLRMs (%). “*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline.
+  Base Model Method All Pair Short Query Pair Short Target Pair
+R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
+  MTomato-Base basic method 71.94 88.22 96.13 44.77 63.77 83.31 45.83 66.08 87.02
+mICL 74.15 90.94 98.26 51.56 75.87 93.84 52.74 78.13 95.82
+late fusion 74.51 91.37 98.45 54.19 78.02 94.52 54.00 79.71 95.80
+NoteLLM-2 74.62* 91.50* 98.47 54.87* 78.82* 94.69* 54.51* 79.75 95.46
+only late fusion 74.23 91.25 98.43 53.20 77.75 94.65 53.65 79.76 95.59
+Omni-Retrieval 71.73 88.12 96.32 44.98 64.26 84.57 45.18 66.14 88.48
+  MQwen-Base basic method 74.02 89.65 97.22 48.15 68.35 88.22 49.82 70.91 91.43
+mICL 75.81 91.57 98.55 53.32 76.40 94.38 54.55 79.32 96.05
+late fusion 75.92 91.50 98.28 52.74 76.03 93.65 54.06 78.89 95.53
+NoteLLM-2 76.49* 92.18* 98.57 55.73* 78.91* 94.94* 56.12* 80.65* 96.24*
+only late fusion 75.84 91.48 98.35 53.33 76.22 93.57 54.27 78.38 95.29
+Omni-Retrieval 74.71 91.05 98.11 51.61 74.34 92.90 51.91 77.07 94.71
+  MQwen-bigG basic method 77.63 92.89 98.91 57.45 80.92 95.73 57.95 83.50 96.99
+mICL 77.96* 93.18 98.92 58.48* 82.09* 96.65* 58.76 84.01 97.20*
+late fusion 77.43 92.93 98.78 57.47 81.78 95.89 58.48 83.99 97.16
+NoteLLM-2 77.56 93.25 98.82 57.94 81.98 96.29 58.76 83.87 97.05
+only late fusion 76.23 92.30 98.67 55.21 79.22 95.58 55.16 81.13 96.70
+Omni-Retrieval 77.38 93.11 98.91 58.26 82.07 96.61 58.28 83.95 97.13
+3. Methodology
 From our investigation, MLRMs trained end-to-end without pre-training are suboptimal, primarily due to the disregard of visual information after processing by LLMs.
@@
 We propose a novel training framework, NoteLLM-2, which contains two methods from different perspectives.
-The first one is from the prompt viewpoint, called mICL.
+The first one is from the prompt viewpoint and is called mICL.
 This method alters the prompt to adjust the attention patterns towards visual information.
@@
 The improved overall framework is shown in Figure 2 (b).
-Specifically, given note $n_{i}$, the mICL mechanism does not attempt compress the multimodal information into one compressed word.
+Specifically, given note $n_{i}$, the mICL does not compress the multimodal information into one compressed word.
 Instead, we separate the multimodal note into two single-modality notes.
-We then adopt a similar ICL way [12, 48] to aggregate the multimodal information.
-We reformulate our note compression prompt as follows: {mdframed} mICL Note Compression Prompt.
-Note content: {’image’: <IMG>}, Compress this note into one word:"<IMG_EMB>".
-Note content: {’title’: $t_{i}$ , ’topic’: $tp_{i}$ , ’content’: $ct_{i}$}.
-Compress this note into one word:" In this template, <IMG_EMB> is a special token.
+We then adopt a similar ICL way (Dong et al., 2022; Wang et al., 2023b) to aggregate the multimodal information.
+We reformulate our note compression prompt as follows:
+{mdframed}
+mICL Note Compression Prompt.
+Note content: {‘image’: ¡IMG¿}, Compress this note into one word:“¡IMG_EMB¿”.
+Note content: {‘title’: $t_{i}$ , ‘topic’: $tp_{i}$ , ‘content’: $ct_{i}$}.
+Compress this note into one word:“ In this template, ¡IMG_EMB¿ is a special token.
 After processing the multimodal embeddings with LLMs, we choose relevant hidden states to represent notes.
-Following [60, 19], we take the previous hidden state of the token <IMG_EMB> as the visual note representation, denoted as $\boldsymbol{n}_{i}^{v}$.
+Following (Zhang et al., 2024a; Jiang et al., 2023a), we take the previous hidden state of the token ¡IMG_EMB¿ as the visual note representation, denoted as $\boldsymbol{n}_{i}^{v}$.
 Due to the causal attention, $\boldsymbol{n}_{i}^{v}$ only contains the note image information.
-We also take the last hidden state as the multimodal note representation, denoted as $\boldsymbol{n}_{i}^{m}$, which can compress visual and textual information.
+We also take the last hidden state as the multimodal note representation, denoted as $\boldsymbol{n}_{i}^{m}$, which contains visual and textual information.
 This method separates the original single embedding into two modality note embeddings.
 The multimodal note representation $\boldsymbol{n}_{i}^{m}$ aggregates useful visual information in a similar manner to ICL.
-The late fusion mechanism adopts original visual embedding to enhance the note embeddings.
-This prevents a textual bias due to the space of LLMs [62] and incorporates more original visual information.
+The late fusion adopts original visual embedding to enhance the note embeddings.
+This prevents a textual bias due to the space of LLMs (Zhang et al., 2024b) and incorporates more original visual information.
 Specifically, after extracting image features by the vision encoder $V_{\theta}$, we take the visual feature, which contains the entire image information, such as the [CLS] token in CLIP ViT-B.
 Then, we use a linear layer to transform these features into the dimension of LLMs, denoted as $\boldsymbol{v}\in\mathbb{R}^{h_{t}}$.
-We adopt the same gate mechanism to fuse the original visual information into the two note embeddings.
-$\boldsymbol{z}$ $=sigmoid(\boldsymbol{W}[\boldsymbol{v},\boldsymbol{n}_{i}^{v}]+\boldsymbol{b}),$ (4)
+We adopt the gate mechanism (Chung et al., 2014) to fuse the original visual information into the note embeddings.
+(7) $\boldsymbol{z}$ $=sigmoid(\boldsymbol{W}[\boldsymbol{v},\boldsymbol{n}_{i}^{v}]+\boldsymbol{b}),$
 $\hat{\boldsymbol{n}_{i}^{v}}$ $=\boldsymbol{z}\odot\boldsymbol{v}+(\boldsymbol{1}-\boldsymbol{z})\odot\boldsymbol{n}_{i}^{v},$
-$\boldsymbol{z}$ $=sigmoid(\boldsymbol{W}[\boldsymbol{v},\boldsymbol{n}_{i}^{m}]+\boldsymbol{b}),$ (5)
-$\hat{\boldsymbol{n}_{i}^{m}}$ $=\boldsymbol{z}\odot\boldsymbol{v}+(\boldsymbol{1}-\boldsymbol{z})\odot\boldsymbol{n}_{i}^{m},$
-where $\hat{\boldsymbol{n}_{i}^{v}}$ and $\hat{\boldsymbol{n}_{i}^{m}}$ means the fused note embeddings.
+where $\hat{\boldsymbol{n}_{i}^{v}}$ means the fused visual note embeddings.
 [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] means the concatenate operation, and $\boldsymbol{W}$ and $\boldsymbol{b}$ are learnable parameters.
 $\odot$ is the element-wise product.
-Next, we adopt the two fused embeddings to conduct contrastive learning as follows:
-$\mathcal{L}^{v}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\hat{\boldsymbol{n}_{i}^{v}},\hat{\boldsymbol{n}_{i}^{v,+}})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\hat{\boldsymbol{n}_{i}^{v}},\hat{\boldsymbol{n}_{j}^{v}})\cdot e^{\tau}}},$ (6)
-$\mathcal{L}^{m}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\hat{\boldsymbol{n}_{i}^{m}},\hat{\boldsymbol{n}_{i}^{m,+}})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\hat{\boldsymbol{n}_{i}^{m}},\hat{\boldsymbol{n}_{j}^{m}})\cdot e^{\tau}}},$ (7)
-where $\mathcal{L}^{v}(\pi,\mu)$ is the loss from visual note embeddings and $\mathcal{L}^{m}(\pi,\mu)$ is the loss from the multimodal note embeddings. The final loss is computed as follows:
-$\mathcal{L}^{f}(\pi,\mu)=\frac{\mathcal{L}^{v}(\pi,\mu)+\alpha\mathcal{L}^{m}(\pi,\mu)}{1+\alpha},$ (8)
+Similarly, we can get the fused multimodal note embedding $\hat{\boldsymbol{n}_{i}^{m}}$ with the same gate operation.
+Next, we adopt the fused note embeddings to conduct contrastive learning as follows:
+(8) $\mathcal{L}^{v}(\pi,\mu)=-\frac{1}{2B}\sum_{i=1}^{2B}log\frac{e^{sim(\hat{\boldsymbol{n}_{i}^{v}},\hat{\boldsymbol{n}_{i}^{v,+}})\cdot e^{\tau}}}{\sum_{j\in[2B]\backslash\{i\}}e^{sim(\hat{\boldsymbol{n}_{i}^{v}},\hat{\boldsymbol{n}_{j}^{v}})\cdot e^{\tau}}},$
+where $\mathcal{L}^{v}(\pi,\mu)$ is the loss from visual note embeddings.
+Similarly, we can get $\mathcal{L}^{m}(\pi,\mu)$, which is the loss from the multimodal note embeddings.
+The final loss is computed as follows:
+(9) $\mathcal{L}^{f}(\pi,\mu)=\frac{\mathcal{L}^{v}(\pi,\mu)+\alpha\mathcal{L}^{m}(\pi,\mu)}{1+\alpha},$
 where $\mathcal{L}^{f}(\pi,\mu)$ is the final loss and $\alpha$ is a hyperparameter.
 In the evaluation, we use $\hat{\boldsymbol{n}_{i}^{m}}$ as the note embeddings, which contain multimodal information.
-4 Experiments
-Table 3: Performance of NoteLLM-2 based on three different MLRMs (%).
-  Method All Pair Short Query Pair Short Target Pair
-R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
-  MTomato-Base 71.94 88.22 96.13 44.77 63.77 83.31 45.83 66.08 87.02
-   + mICL 74.16 90.95 98.26 51.56 75.86 93.84 52.71 78.12 95.82
-   + late fusion 74.51 91.37 98.45 54.19 78.02 94.50 54.00 79.71 95.80
-   + NoteLLM-2 74.61 91.50 98.47 54.87 78.81 94.69 54.52 79.75 95.46
-only late fusion 74.22 91.25 98.43 53.21 77.75 94.65 53.65 79.77 95.59
-  MQwen-Base 74.02 89.65 97.22 48.15 68.35 88.22 49.82 70.91 91.43
-   + mICL 75.82 91.57 98.55 53.32 76.40 94.38 54.52 79.31 96.05
-   + late fusion 75.92 91.49 98.28 52.74 76.03 93.65 54.06 78.89 95.53
-   + NoteLLM-2 76.50 92.18 98.58 55.72 78.91 94.94 56.13 80.65 96.24
-only late fusion 75.84 91.49 98.35 53.36 76.23 93.57 54.27 78.39 95.30
-  MQwen-bigG 77.64 92.89 98.91 57.45 80.92 95.73 57.90 83.49 96.99
-   + mICL 77.97 93.19 98.92 58.50 82.08 96.66 58.58 84.01 97.20
-   + late fusion 77.43 92.92 98.78 57.47 81.77 95.89 58.50 83.99 97.16
-   + NoteLLM-2 77.56 93.38 98.82 57.94 81.98 96.29 58.79 83.87 97.05
-only late fusion 76.24 92.31 98.67 55.21 79.22 95.58 55.15 81.13 96.70
-4.1 Performance Evaluation
-We conduct experiments on all components of NoteLLM-2 based on three MLRMs to show our method effectiveness.
-Besides, we perform an ablation study, only late fusion, which only use late fusion to integrate image and text information, without inputting image embeddings into LLMs.
+4. Experiments
+4.1. Performance Evaluation
+We conduct experiments on following baselines based on three MLRMs to show NoteLLM-2’s effectiveness.
+basic method is the naive representation method of MLRMs.
+mICL adds the mICL method based on the basic method.
+late fusion integrates the late fusion mechanism into the basic method.
+NoteLLM-2 contains both mICL and late fusion simultaneously.
+only late fusion only uses late fusion to integrate image and text information, without inputting image embeddings into LLMs.
+Omni-Retrieval (Yu et al., 2022a) adopts cross-modal contrastive learning across various embeddings, including image-only, text-only, and multimodal, to achieve alignment.
 The results are presented in Table 4.
-Firstly, NoteLLM-2 can significantly enhance the overall performance for MTomato-Base and MQwen-Base, which have relatively small vision encoders.
+Firstly, NoteLLM-2 can significantly enhance the overall performance of MTomato-Base and MQwen-Base, which have relatively small vision encoders.
 Additionally, NoteLLM-2 primarily enhances MQwen-bigG in terms of short pairs.
 Secondly, mICL enhances the performance of all models, while late fusion is more effective in models with relatively small vision encoders.
-Concurrently, we find that only late fusion is a straightforward and efficient fusion method.
-However, its simplicity could potentially be a drawback when the vision encoder is more powerful, as it may not fully and effectively interact with LLMs, leading to a loss in performance.
-Lastly, MQwen-bigG + NoteLLM-2 still underperforms compared to Qwen-VL-Chat, especially in shorter pairs.
-We attribute this to the significant difference in visual token length between MQwen-bigG ($16$) and Qwen-VL-Chat ($256$).
-((a)) MTomato-Base
-((b)) MQwen-Base
-((c)) MQwen-bigG
-Figure 4: Relative sizes of saliency scores in different layers of different MLRMs.
-4.2 Saliency Scores of Enhanced MLRMs
+When the vision encoders are powerful, MQwen-bigG+late fusion improves performance on short pairs by focusing more on images.
+However, this overemphasis on images can detrimentally affect text comprehension, reducing overall performance and weakening the mICL method that utilizes LLM prompts.
+We will design a more balanced late fusion mechanism in future work.
+Concurrently, only late fusion is a straightforward and efficient fusion method.
+However, its simplicity could potentially be a drawback when the vision encoder is more powerful, as it may not fully interact with LLMs, leading to a loss in performance.
+Besides, the enhancement of Omni-Retrieval for visual information is still limited compared to NoteLLM-2.
+This is because Omni-Retrieval ignores the distinct mechanisms of LLMs, and other cross-modal losses may interfere with multimodal embeddings.
+Lastly, MQwen-bigG+NoteLLM-2 underperforms compared to Qwen-VL-Chat, especially in shorter pairs.
+This is due to the visual token length difference between MQwen-bigG ($16$) and Qwen-VL-Chat ($256$).
+4.2. Saliency Scores of Enhanced MLRMs
 To further understand the impact of NoteLLM-2 on MLRMs, we illustrate the difference in saliency scores of enhanced MLRMs.
-The saliency scores for the original fine-tuned method are $S_{v}$, $S_{t}$ and $S_{o}$.
+The saliency scores for the original fine-tuned method are $S_{v}$, $S_{t}$, and $S_{o}$.
 The saliency scores for the enhanced fine-tuned method are $\hat{S_{v}}$, $\hat{S_{t}}$ and $\hat{S_{o}}$.
 We consider the visual note compressed word as a part of the vision embeddings $E_{v}$.
-Our results are shown in Figure 4.
-The representations of all enhanced MLRMs amplify the direct focus on images while reducing the emphasis on $S_{o}$ in the shallow layers.
+Our results are shown in Figure LABEL:fig:ssdif.
+The representations of all enhanced MLRMs amplify $S_{v}$, the direct focus on images, while reducing the emphasis on $S_{o}$ in the shallow layers.
 At the same time, $S_{t}$ remains virtually unchanged.
-This is attributed to mICL’s use of the same compressed prompt to compress both modalities.
-By identifying similar compression patterns from image information, mICL can bolster the concentration of multimodal representations on images, similar to ICL [48].
-Table 4: Impact of the number of visual tokens $L$ (%).
-  #Visual Tokens All Pair Short Query Pair Short Target Pair
+More balanced $S_{v}$ and $S_{o}$ mean that the information flow gain brought by introducing images not only highlights the relevant multimodal embeddings but is also aggregated by the final compressed embedding directly.
+This is helpful for final compressed embeddings to capture crucial patterns in images that are totally different from text.
+This result is attributed to mICL’s use of the same compressed prompt to compress both modalities.
+By identifying similar compression patterns from images, mICL can bolster the concentration of multimodal representations on images, similar to ICL (Wang et al., 2023b).
+4.3. Hyper-Parameter Analysis
+In this section, we use MTomato-Base+NoteLLM-2 to conduct hyper-parameter analysis experiments.
+Impact of length of visual tokens.
+We explore the impact of visual token length.
+The results are shown in Table 5.
+Reducing the length from $16$ to $8$ leads to poorer performance, while inference is more efficient.
+There exists a tradeoff between performance and inference speed concerning the length of visual tokens.
+Moreover, MTomato-Base+NoteLLM-2 with 8 visual tokens outperforms MTomato-Base with 16 visual tokens in Table 2.
+This shows that our method enhances visual attention by modified attention patterns and fusion, rather than merely expanding image token space.
+Impact of text and visual loss ratio.
+We vary the ratio of text to visual loss $\alpha$, as shown in Table 6.
+When the ratio is small, the performance is slightly worse.
+However, as we increase the ratio, we find that our method is insensitive to changes in the text to visual loss ratio.
+Table 5. Impact of the number of visual tokens $L_{c}$ (%). Inf. Sp. means the inference speed.
+  $L_{c}$ Inf. Sp. All Pair Short Query Pair Short Target Pair
 R@100 R@1k R@10k R@100 R@1k R@10k R@100 R@1k R@10k
-  8 74.17 90.85 98.16 52.80 76.13 93.76 53.19 77.87 94.48
-16 74.61 91.50 98.47 54.87 78.81 94.69 54.52 79.75 95.46
-32 74.55 91.41 98.38 54.56 78.91 94.94 54.21 79.83 95.63
-48 74.30 91.57 98.48 54.31 78.83 95.13 53.73 80.06 95.92
-Table 5: Impact of textual loss and visual loss ratio $\alpha$ (%).
+  8 51.0 74.17 90.85 98.16 52.80 76.13 93.76 53.19 77.87 94.48
+16 48.2 74.62 91.50 98.47 54.87 78.82 94.69 54.51 79.75 95.46
+32 44.2 74.55 91.41 98.38 54.56 78.91 94.94 54.21 79.83 95.63
+48 41.0 74.30 91.57 98.48 54.31 78.83 95.13 53.73 80.06 95.92
+Table 6. Impact of textual loss and visual loss ratio $\alpha$ (%).
   $\alpha$ All Pair Short Query Pair Short Target Pair
@@
 3 74.38 91.41 98.42 54.15 77.71 94.59 54.06 79.31 95.53
-9 74.61 91.50 98.47 54.87 78.81 94.69 54.52 79.75 95.46
+9 74.62 91.50 98.47 54.87 78.82 94.69 54.51 79.75 95.46
 19 74.53 91.33 98.54 54.56 78.68 95.00 53.92 80.04 95.53
-4.3 Hyper-Parameter Analysis
-In this section, we use MTomato-Base to conduct hyper-parameter analysis experiments.
-Impact of length of visual tokens.
-We explore the impact of visual token length.
-The results are shown in Table 4.
-Reducing the length from $16$ to $8$ worsens performance in short pairs.
-There exists a tradeoff between performance and inference speed concerning the length of visual tokens.
-Impact of text and visual loss ratio.
-We vary the ratio of text to visual loss $\alpha$, as shown in Table 5.
-When the ratio is small, the performance is slightly worse.
-However, as we increase the ratio, we find that our method is insensitive to changes in the text to visual loss ratio.
-5 Conclusion
-This paper explores using LLMs to improve textual comprehension in multimodal representation tasks under I2I recommendation scenarios.
-We design an end-to-end training method that can customize the integration of any existing LLMs and vision encoders.
-This approach reduces reliance on open-source MLLMs, which require costly multimodal pre-training.
-To prevent the problem of ignoring visual information in end-to-end training, we propose NoteLLM-2, which contains mICL and late fusion.
-The effectiveness of this approach is confirmed through extensive experiments.
-Appendix
-Appendix A Related Work
-A.1 Multimodal Representation
+Figure 9.
+The online serving pipeline includes offline and online phases.
+The offline phase generates embeddings with the MLRM trained with NoteLLM-2 and updates the ANN index with new notes.
+The online phase recalls notes from the index based on the user’s interaction history.
+Table 7.
+Comparison of different LLMs and MLRMs for representation tasks.
+$L_{c}$ means the number of vision embeddings for each image.
+All models are trained on A100 GPUs.
+The training cost is measured in GPU hours (GPU-h).
+The inference speed test is conducted on an A100 GPU, with an inference batch size set at $50$.
+All notes in this test are the longest note in the dataset.
+The speed is quantified as the number of longest notes inferred per second.
+  Method Input Vision Encoder LLM VL Connector # Total Par. $L_{c}$ Training GPU-h Inf. Speed
+  Tomato Text - Tomato (6.7B) - 6.7B - 78.3 59.0
+Qwen-Chat Text - Qwen-Chat (7.1B) - 7.1B - 86.0 57.0
+  BLIP-2 Multimodal ViT-g (1.0B) OPT (6.7B) Q-Former (105.1M) 7.8B 32 110.8 44.2
+Qwen-VL-Chat Multimodal ViT-bigG (1.8B) Qwen-Chat (7.1B) Resamplers (76.1M) 9.0B 256 432.6 11.5
+  MTomato-Base Multimodal ViT-B (85.8M) Tomato (6.7B) Q-Former (49.6M) 6.8B 16 108.6 51.5
+    + NoteLLM-2 Multimodal ViT-B (85.8M) Tomato (6.7B) Q-Former (49.6M) 6.9B 16 110.7 48.2
+  MQwen-Base Multimodal ViT-B (85.8M) Qwen-Chat (7.1B) Q-Former (49.6M) 7.3B 16 100.3 50.5
+    + NoteLLM-2 Multimodal ViT-B (85.8M) Qwen-Chat (7.1B) Q-Former (49.6M) 7.3B 16 101.6 46.9
+  MQwen-bigG Multimodal ViT-bigG (1.8B) Qwen-Chat (7.1B) Q-Former (142.9M) 9.1B 16 129.8 36.8
+    + NoteLLM-2 Multimodal ViT-bigG (1.8B) Qwen-Chat (7.1B) Q-Former (142.9M) 9.2B 16 123.1 35.2
+4.4. Efficiency Analysis
+We provide a comparison of the model parameter size, training cost, and inference speed of LLMs in Table 7.
+State-of-the-art MLLMs such as Qwen-VL-Chat, designed for multimodal understanding and answering, have high training costs for adaptation and slow inference speed, making them impractical for fast-paced recommendation tasks.
+On the other hand, MLRMs designed for representation tasks are more efficient for both training and inference.
+Additionally, due to NoteLLM-2’s longer prompt template, integrating NoteLLM-2 usually slightly slows down training and inference, but achieves better performance.
+4.5. Online Serving and Experiments
+We present our online serving pipeline in Figure 9, which consists of offline and online phases.
+During the offline phase, the MLRM trained with NoteLLM-2 generates multimodal embeddings in real-time for newly generated notes from users.
+These embeddings are saved in an embedding table for reuse.
+Furthermore, the I2I ANN index is dynamically updated with embeddings of new notes.
+In the online phase, the system selects history embeddings from the existing embedding table based on the user’s interaction history notes.
+Using these history note embeddings, the system performs multiple ANN searches to find relevant notes in the I2I index.
+The top-k relevant notes are recalled and passed to the next recommendation stage.
+Now, our model has replaced the previous I2I recall process and has already become a crucial recall channel on the platform.
+We conduct week-long online experiments on our platform by allocating 10% of the overall traffic for A/B testing.
+The online baseline uses the same LLM as our online method for I2I training, with the training approach utilizing NoteLLM (Zhang et al., 2024a).
+Our MLRM trained with NoteLLM-2 enhances the click count of the first thousand exposures by $6.35\%$ and the number of interactions within the first $24$ hours of note publication by $8.08\%$.
+Besides, notes with the first interaction within 24 hours and 100 exposures increase by $9.83\%$.
+5. Related Work
+Multimodal Representation.
 With the advancement of pre-training techniques, numerous studies utilize web-scale image-text data to pre-train models for a fundamental understanding of multimodal information.
 There are two main pre-training techniques.
-The first technique adopts generative pre-training objectives, following masked language/image modeling [14, 49, 27] or image captioning [56, 26].
-The second technique uses contrastive pre-training objectives, such as contrastive learning [37, 58, 45, 46, 56, 49] and image-text matching tasks [14, 27, 26].
+The first technique adopts generative pre-training objectives, following masked language/image modeling (Dou et al., 2022; Wang et al., 2023a; Li et al., 2021) or image captioning (Yu et al., 2022b; Li et al., 2022).
+The second technique uses contrastive pre-training objectives, such as contrastive learning (Radford et al., 2021; Zhai et al., 2023; Sun et al., 2023, 2024; Yu et al., 2022b; Wang et al., 2023a) and image-text matching tasks (Dou et al., 2022; Li et al., 2021, 2022).
 These models demonstrate a powerful zero-shot ability in many downstream tasks.
-However, the text comprehension capabilities of these models are unsatisfactory [6, 59], particularly in the era of LLMs.
-In this work, we explore to use LLMs to enhance the multimodal representations.
-A.2 Large Language Models
-LLMs have attracted extensive attention due to their remarkable text understanding and generation capability [18, 7, 2, 47].
-Therefore, numerous text applications have been developed based on LLMs [31, 60, 34, 52, 22].
-To understand more modality information, LLMs are extended into MLLMs [3, 5, 55, 33, 24].
-Most MLLMs employ specialized modality encoders to capture non-textual information [42, 3, 5].
-This information is then mapped into textual LLM space using connectors and subsequently processed by LLMs.
-Due to the significant understanding ability of LLMs, MLLMs can generate answers according to textual instruction and multimodal information [15, 13, 3, 5].
-While existing studies primarily focus on the generation abilities of LLMs, some explore the text representation ability of LLMs [34, 41, 60, 19, 31, 4], demonstrating significant potential to enhance performance for numerous downstream tasks.
-However, no studies have explored the multimodal representation ability with the assistance of LLMs.
-Therefore, we aim to investigate the multimodal representation ability enhanced by LLMs in recommendation scenarios.
-A.3 Multimodal I2I Recommendation
-I2I recommendation generates a ranked list of items from a vast pool based on a specific target item, which is an important technique for recall stages in recommendation [54, 64, 60].
-I2I recommendation employs either a pre-constructed I2I index [54] or the approximate k-nearest neighbor method [21] to retrieve relevant items online.
-Some studies have investigated multimodal content-based I2I recommendations [57, 10, 20, 63, 16].
-For example, CommerceMM [57] adopts multimodal pre-training tasks used in METER [14] and omni-retrieval tasks to train commerce multimodal representation.
-MAPS [10] proposes a multimodal attention framework to measure product similarity.
+However, the text comprehension capabilities of these models are unsatisfactory (Chen et al., 2023; Zhang et al., 2024c), particularly in the era of LLMs.
+In this work, we explore the use of LLMs’ powerful comprehension capabilities to enhance multimodal representations.
+Large Language Models.
+LLMs have attracted extensive attention due to their remarkable text understanding and generation capability (Jiang et al., 2023b; Chiang et al., 2023; Bai et al., 2023a; Touvron et al., 2023).
+Therefore, numerous text applications have been developed based on LLMs (Ma et al., 2023; Zhang et al., 2024a; Muennighoff et al., 2024; Xu et al., 2024; Karamcheti et al., 2024).
+To understand more modality information, LLMs are extended into MLLMs (Bai et al., 2023b; Chen et al., 2024; Yin et al., 2023; McKinzie et al., 2024; Laurençon et al., 2024).
+Most MLLMs employ specialized modality encoders to capture non-textual information (Su et al., 2023; Bai et al., 2023b; Chen et al., 2024).
+This information is mapped into textual space and processed by LLMs, enabling MLLMs to generate answers from multimodal input (Ge et al., 2023; Dong et al., 2023; Bai et al., 2023b; Chen et al., 2024).
+While existing studies primarily focus on the generation abilities of LLMs, some explore the text representation ability of LLMs (Muennighoff et al., 2024; Springer et al., 2024; Zhang et al., 2024a; Jiang et al., 2023a; Ma et al., 2023; BehnamGhader et al., 2024).
+However, no studies have explored multimodal representation with LLMs.
+Therefore, we investigate the multimodal representation ability enhanced by LLMs in recommendation scenarios.
+Multimodal I2I Recommendation.
+I2I recommendation generates a ranked list of items from a vast pool based on item similarity, which is an important technique for recall stages in recommendation (Yang et al., 2020; Zhu et al., 2018; Zhang et al., 2024a).
+I2I recommendation employs a pre-constructed I2I index (Yang et al., 2020) to retrieve relevant items by the approximate k-nearest neighbor method (Johnson et al., 2019).
+Some studies have investigated multimodal content-based I2I recommendations (Yu et al., 2022a; Das et al., 2022; Jin et al., 2023; Zheng et al., 2023; Han et al., 2023).
 However, these works adopt small-scale pre-trained visual-linguistic models as backbones, which have huge potential for scaling up.
-As shown in [60, 19, 4], LLMs can outperform smaller models in textual representation tasks significantly.
-Therefore, in this work, we explore empowering the multimodal representations with LLMs in multimodal I2I recommendation.
-Appendix B Limitations and Future Works
-More modalities.
-In this work, our focus is primarily on the processing of multimodal situations that involve the input of a single image and corresponding text information.
-This is the most simplified situation within our research scenarios.
-In fact, the scope of our representation method can extend to include audio and video data [42, 32, 28].
-This direction presents an opportunity for further exploration and expansion into more complex modalities.
-General scenarios.
-Our MLRMs mainly focus on the multimodal representation ability in recommendation scenarios.
-These scenarios determine item similarity by analyzing the interaction between items and users.
-However, this is not a universally applicable scenario.
-The current focus of our MLRMs is somewhat specialized and may not cover all potential use cases.
-Given the potential of MLRMs, it is promising to consider broadening their multimodal representation ability into more general scenarios [35].
-Larger and stronger base models.
-Due to the rapid development of LLMs and vision encoders, there are an extensive range of larger and more powerful base models available in the market [58, 2, 11].
-However, due to resource limitations and the demand for rapid recommendations, we have not yet experimented with all of these models.
-There is still a vast scope to investigate more advanced models, which could potentially boost performance even further.
-Appendix C Dataset Details
-Table 6: Detailed statistics of training and testing dataset.
-                                  training dataset
-  # notes 1,487,613 # note pairs 1,130,906
-avg. # words per title 12.38 avg. # topic per note 5.26
-avg. # words per topic 4.53 avg. # words per content 133.20
-                                  testing dataset
-  # notes 534,144 # note pairs 21,262
-avg. # words per title 11.68 avg. # topic per note 4.58
-avg. # words per topic 4.50 avg. # words per content 116.00
-Appendix D Training Details
-Table 7: Detailed training hyperparameters for training MLRMs.
+Therefore, in this work, we explore empowering the multimodal representations with LLMs in multimodal I2I recommendations.
+6. Conclusion and Future Work
+This paper explores using LLMs to improve textual comprehension in multimodal representation tasks under I2I recommendation scenarios.
+We design an end-to-end fine-tuning method that can customize the integration of any existing LLMs and vision encoders.
+This approach reduces reliance on open-source MLLMs, which require costly multimodal pre-training.
+To prevent the problem of ignoring visual information in end-to-end fine-tuning, we propose NoteLLM-2, which contains mICL and late fusion.
+The effectiveness of this approach is confirmed through extensive experiments.
+In future work, we will extend our method to more complex modalities, like videos, and broader scenarios and recommendation tasks.
+Acknowledgements
+This work was supported in part by the grants from National Science and Technology Major Project (No.
+2023ZD0121104), and National Natural Science Foundation of China (No.62222213, 62072423).
+Xiangyu Zhao was partially supported by Research Impact Fund (No.R1015-23), APRC - CityU New Research Initiatives (No.9610565, Start-up Grant for New Faculty of CityU), CityU - HKIDS Early Career Research Grant (No.9360163), Hong Kong ITC Innovation and Technology Fund Midstream Research Programme for Universities Project (No.ITS/034/22MS), Hong Kong Environmental and Conservation Fund (No.
+88/2022), SIRG - CityU Strategic Interdisciplinary Research Grant (No.7020046), Huawei (Huawei Innovation Research Program), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Ant Group (CCF-Ant Research Fund, Ant Group Research Fund), Alibaba (CCF-Alimama Tech Kangaroo Fund No.
+2024002), CCF-BaiChuan-Ebtech Foundation Model Fund, Collaborative Research Fund No.C1043-24GF, and Kuaishou.
+Appendix
+Appendix A Training Details
+Table 8. Detailed training hyperparameters for fine-tuning MLRMs.
   Configuration Hyperparameter
@@
 Training steps $7,956$
-For all models with LLMs, we list the detailed training hyperparameter in Table 7. Besides, we use DeepSpeed [38], and Zero Redundancy Optimizer (ZeRO) [39] Stage 3 to train our models.
-Appendix E Model Details
-E.1 Baseline Details
-To evaluate the effectiveness of MLRMs, we compare them with several baselines.
-(1) CLIP ViT-B [37] is trained using only image input.
-(2) BM25 [40] is a basic baseline, which counts the relevance between the query and documents based on every word in the query.
-(3) RoBERTa-wwm-ext [9] is a classic traditional Chinese BERT-like [23] text encoder with pure text input.
-(4) Tomato is our continually pre-trained LLM that is based on LLaMA 2 [47] and utilizes data from our platform for training.
-(5) NoteLLM [60] unifies the embedding generation and topic generation into one task.
-(6) METER Merge-attn and METER Co-attn [14] leverage best training strategy to train vision-and-language transformers.
-METER Merge-attn adopts the merged attention as the fusion module, while METER Co-attn uses co-attention as the fusion module.
-All baselines are trained on our training dataset.
-E.2 End-to-End MLRM Details
-We provide the details in Table 8.
-Table 8: Model settings of end-to-end MLRMs.
+For all models with LLMs, we list the detailed training hyperparameter in Table 8.
+Besides, we use DeepSpeed (Rasley et al., 2020), and Zero Redundancy Optimizer (ZeRO) (Ren et al., 2021) Stage 3 to train our models.
+Appendix B End-to-End MLRM Details
+We provide the details about end-to-end MLRMs in Table 9.
+Besides, we present the continuous pre-training details of Tomato to enhance understanding of its characteristics.
+Tomato is based on LLaMA 2 (Touvron et al., 2023), which originally only supports English.
+However, since our platform targets the Chinese market, it is necessary to enhance LLaMA 2’s Chinese understanding capabilities.
+Thus, we have made the following improvements to continuous pre-training:
+  • Vocabulary Expansion: We extended the original vocabulary from $32,000$ tokens to $49,216$ tokens. The additional $17,216$ tokens are primarily Chinese.
+  • Pre-training Data: We collect pre-training data from three sources: Chinese corpora, English corpora, and Platform-specific Data.
+We utilize open-source Chinese corpora, including Wanjuan (He et al., 2023) and WuDao (Xue et al., 2022), totaling approximately 2T tokens.
+To maintain English proficiency, we also include RedPajama (Computer, 2023).
+Besides, we incorporate 30B tokens of high-quality notes collected from our platform to better understand the data characteristics of our platform.
+Tomato enhances Chinese understanding while maintaining its English capabilities at a level comparable to the original LLaMA 2 model.
+It is worth noting that the pre-training details are unrelated to our contributions.
+Our method is based on the model after continual pre-training.
+This also shows an advantage of our method, which allows for the personalized enhancement of the base models for different modalities before fine-tuning using NoteLLM-2.
+Table 9. Model settings of end-to-end MLRMs.
   Configuration MTomato-Base MQwen-Base MQwen-bigG
@@
 LLM intermediate size $11,008$ $22,016$ $22,016$
-E.3 Comparison of MLRMs
-We provide the comparison in Table 9.
-Table 9: Comparison of different LLMs for representation tasks.
-$L_{c}$ means the number of vision embeddings for each image.
-The inference speed test is conducted on an A100 GPU, with an inference batch size set at $50$.
-All notes in this test are the longest note in the dataset.
-The speed is quantified as the number of longest notes inferred per second.
-  Method Vision Encoder LLM VL Connector # Total Par. $L_{c}$ Inf. Speed
-  Tomato - Tomato (6.7B) - 6.7B - 59.0
-Qwen-Chat - Qwen-Chat (7.1B) - 7.1B - 57.0
-  BLIP-2 ViT-g (1.0B) OPT (6.7B) Q-Former (105.1M) 7.8B 32 44.2
-LLaVA-1.5 ViT-L (303.5M) Vicuna (6.7B) MLP (21.0M) 7.1B 576 12.1
-Qwen-VL-Chat ViT-bigG (1.8B) Qwen-Chat (7.1B) Resamplers (76.1M) 9.0B 256 11.5
-MTomato-Base ViT-B (85.8M) Tomato (6.7B) Q-Former (49.6M) 6.8B 16 51.5
-    + NoteLLM-2 ViT-B (85.8M) Tomato (6.7B) Q-Former (49.6M) 6.9B 16 48.2
-MQwen-Base ViT-B (85.8M) Qwen-Chat (7.1B) Q-Former (49.6M) 7.3B 16 50.5
-    + NoteLLM-2 ViT-B (85.8M) Qwen-Chat (7.1B) Q-Former (49.6M) 7.3B 16 46.9
-MQwen-bigG ViT-bigG (1.8B) Qwen-Chat (7.1B) Q-Former (142.9M) 9.1B 16 36.8
-    + NoteLLM-2 ViT-bigG (1.8B) Qwen-Chat (7.1B) Q-Former (142.9M) 9.2B 16 35.2
-Appendix F More Information about Saliency Scores
-F.1 Formal Definition of $S_{v}$, $S_{t}$ and $S_{o}$
-$S_{v}$, the mean significance of information flow from the vision embeddings $E_{v}$ to the final compressed position $c$:
-$S_{v}$ $=\frac{\sum_{(i,j)\in P_{v}}I_{l}(i,j)}{\left|P_{v}\right|},$ (9)
-$P_{v}$ =$=\{(c,j):j\in E_{v}\}.$
-$S_{t}$, the mean significance of information flow from the textual embeddings $E_{t}$ to the final compressed position $c$:
-$S_{t}$ $=\frac{\sum_{(i,j)\in P_{t}}I_{l}(i,j)}{\left|P_{t}\right|},$ (10)
-$P_{t}$ =$=\{(c,j):j\in E_{t},j<c\}.$
-$S_{o}$, the mean significance of information flow amongst all words, excluding influences represented by $S_{v}$ and $S_{t}$:
-$S_{o}$ $=\frac{\sum_{(i,j)\in P_{o}}I_{l}(i,j)}{\left|P_{o}\right|},$ (11)
-$P_{o}$ =$=\{(i,j):j<i\}-P_{v}-P_{t}.$
-F.2 Results of Qwen-VL-Chat
-Figure 5: Relative sizes of $S_{v}$, $S_{t}$ and $S_{o}$ in different layers of Qwen-VL-Chat.
-We observe a unique pattern of saliency scores in Qwen-VL-Chat, as shown in Figure 5.
-The predominant information flow in Qwen-VL-Chat is textual, with a value higher than $0.8$ in all layers.
-This result is attributed to the extensive visual embeddings ($256$ in Qwen-VL-Chat) inputted into LLMs, which reduces the average saliency of image information flow.
-This denotes that traditional MLLMs are not directly suitable for MLRMs.
-The necessary action to adapt MLLMs to MLRMs is enhancing the efficiency of information flow.
-Appendix G Performance of Zero-shot Multimodal Representation
-Table 10: Zero-shot representation performance of different models (%).
+Table 10.
+Zero-shot representation performance of different models under different modal inputs (%).
+“*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline.
   Method Input All Pair Short Query Pair Short Target Pair
@@
 Text 30.86 46.71 64.14 21.90 34.43 49.63 22.28 32.98 47.48
-Multimodal 37.01 55.46 73.71 30.51 47.78 68.73 31.56 48.09 68.13
+Multimodal 37.01 55.45 73.72 30.50 47.78 68.74* 31.53 48.09 68.12
   Qwen-VL Image 17.55 31.62 51.58 17.77 33.55 55.56 18.41 33.88 55.97
 Text 38.74 54.65 70.37 23.31 35.17 51.95 22.17 34.84 51.29
-Multimodal 37.88 55.51 74.73 26.28 43.20 64.89 29.82 47.63 69.72
+Multimodal 37.87 55.48 74.72 26.29 43.14 64.88 29.81 47.66 69.68*
   Qwen-VL-Chat Image 17.43 31.15 50.48 17.59 32.29 55.04 18.75 33.40 55.09
 Text 37.92 53.02 68.48 22.63 33.87 49.15 21.32 32.50 48.03
-Multimodal 40.93 58.59 75.44 29.10 44.48 64.14 31.81 49.09 68.57
+Multimodal 40.94 58.58 75.43 29.07 44.51 64.14 31.81 49.10 68.54
   Qwen-Chat Text 38.98 54.08 69.02 23.08 34.90 50.44 26.04 38.35 53.33
 Tomato Text 42.67 60.54 78.01 24.51 37.35 56.45 32.27 46.85 66.75
-BM25 Text 55.20 72.67 82.83 35.82 48.42 59.71 35.07 49.63 63.68
+BM25 Text 55.22* 72.64* 82.82* 35.75* 48.41* 59.81 35.05* 49.52* 63.64
+Appendix C Saliency Scores of Qwen-VL-Chat
+We observe a unique pattern of saliency scores in Qwen-VL-Chat, as shown in Figure 10.
+The predominant information flow in Qwen-VL-Chat is textual, with a value higher than $0.8$ in all layers.
+This result is attributed to the extensive visual embeddings ($256$ in Qwen-VL-Chat) inputted into LLMs, which reduces the average saliency of image information flow.
+This denotes that traditional MLLMs are not directly suitable for MLRMs.
+The necessary action to adapt MLLMs to MLRMs is to increase the density of information flow.
+Figure 10. Relative sizes of $S_{v}$, $S_{t}$ and $S_{o}$ in different layers of Qwen-VL-Chat.
+Appendix D Performance of Zero-shot Multimodal Representation
 In this section, we explore the multimodal representation ability of MLLMs in a zero-shot way in the multimodal I2I recommendation task.
-We select four popular MLLMs, i.e., BLIP-2 [25], LLaVA-1.5 [29], Qwen-VL [3] and Qwen-VL-Chat [3].
-We also compare some baselines BM25 [40], Qwen-Chat [2] and our continual pre-trained LLM Tomato, which is pre-trained on our platform data based on LLaMA 2 [47], but lacks vision perception ability.
+We select four popular MLLMs, i.e., BLIP-2 (Li et al., 2023), LLaVA-1.5 (Liu et al., 2023b), Qwen-VL (Bai et al., 2023b) and Qwen-VL-Chat (Bai et al., 2023b).
+We also compare some baselines BM25 (Robertson et al., 2009), Qwen-Chat (Bai et al., 2023a) and our continual pre-trained LLM Tomato, which is pre-trained on our platform data based on LLaMA 2 (Touvron et al., 2023), but lacks vision perception ability.
 The MLLMs details are concluded in Table 7.
@@
 Lastly, we find that LLMs can achieve performance comparable to MLLMs without any image input.
-Appendix H Case Study
-Figure 6: The visualization cases of some methods.
-Each line is a case, which the left two columns are the query note and target note, the other columns are the top-$1$ ranked notes for each method.
-We also show the target note ranking.
-We show some cases for multimodal I2I note recommendation as shown in Figure 6.
-In the first case, both the query note and the target note discuss a guide to Oahu, an island in Hawaii.
-MTomato-Base + NoteLLM-2 successfully ranks the target note in the top position.
-While METER Co-attn and MTomato-Base also recall relevant notes in the first place, they differ slightly.
-One talks about an unpleasant experience in Oahu, and the other discusses an overall trip plan for Hawaii.
-The second example illustrates a note pair about the Pagani car model.
-Although all baselines recall notes about various car models, these notes discuss many different types of cars.
-MTomato-Base + NoteLLM-2 ranks the Pagani model car in the first place, which can be attributed to both the image and the text content.
-The third case discusses how to make good coffee and coffee tricks.
-LLM-based methods recall relevant notes about the tools needed to make coffee and how to do it.
-In contrast, METER Co-attn, which has a weaker understanding of text, only recommends coffee beans.
-The last case involves a short note pair.
-The query note is about an AI drawing for China’s National Day, and the target note is about a mooncake drawn by AI, which relates to the Chinese traditional festival, the Mid-Autumn Festival.
-We find that MTomato-Base + NoteLLM-2 can understand the query note image about Chinese culture and recommends notes talking about Chang’e, a figure from Chinese folklore associated with the Mid-Autumn Festival.
-Tomato and MTomato-Base only focus on the AI painting, rather than the image content.
+Appendix E Model Performance on Long Query and Target Pairs
+To comprehensively evaluate our method’s effectiveness, we conduct tests on Long Query and Long Target Pairs.
+Long notes are defined as those exceeding $165$ tokens (about $10\%$ of test notes).
+The test dataset contains 2,228 long query pairs and 2,177 long target pairs.
+The results are shown in Table 11.
+Due to the excellent long-context understanding of LLMs, the performance on long pairs is significantly higher than on overall pairs, as long pairs provide more information.
+Additionally, our method improves upon the basic method for long pairs, which shows that enhancing the focus on visual information is also important for long pairs.
+Table 11. Performance of different methods based on three MLRMs on Long Query Pair and Long Target Pair (%).
+  Base Model Method Long Query Pair Long Target Pair
+R@100 R@1k R@10k R@100 R@1k R@10k
+  MTomato-Base basic method 80.41 95.06 98.93 78.21 94.07 98.56
+mICL 80.70 95.50 99.37 79.64 94.12 98.81
+late fusion 80.85 95.79 99.41 79.84 94.36 99.01
+NoteLLM-2 81.57 95.55 99.32 79.34 95.05 99.25
+only late fusion 80.80 95.30 99.32 80.23 94.76 98.91
+Omni-Retrieval 79.98 94.68 98.98 77.77 93.13 98.67
+  MQwen-Base basic method 80.80 95.02 99.23 79.99 94.47 98.86
+mICL 82.01 95.31 99.37 80.98 94.37 99.01
+late fusion 82.16 95.50 99.42 81.07 94.51 99.11
+NoteLLM-2 82.01 95.99 99.47 81.47 95.01 99.31
+only late fusion 81.96 95.93 99.46 82.11 94.56 99.06
+Omni-Retrieval 80.95 95.55 99.23 79.25 94.32 98.72
+  MQwen-bigG basic method 82.78 96.17 99.60 82.31 95.75 99.01
+mICL 83.36 96.17 99.61 82.51 95.80 99.16
+late fusion 82.65 95.93 99.56 82.26 95.16 99.21
+NoteLLM-2 82.74 96.32 99.56 81.82 95.80 99.01
+only late fusion 82.64 95.79 99.46 80.78 95.35 99.25
+Omni-Retrieval 82.30 96.03 99.41 80.63 95.16 99.11